### v1

# Beyond the Request: Harnessing HTTP Response Headers for Cross-Browser Web Tracker Classification in an Imbalanced Setting

Abstract.
The World Wide Web’s connectivity is greatly attributed to the HTTP protocol, with HTTP messages offering informative header fields that appeal to disciplines like web security and privacy, especially concerning web tracking. Despite existing research employing HTTP/S request messages to identify web trackers, HTTP/S response headers are often overlooked. This study endeavors to design effective machine learning classifiers for web tracker detection using HTTP/S response headers. Data from the Chrome, Firefox, and Brave browsers, obtained through the traffic monitoring browser extension T.EX, serves as our data set. Eleven supervised models were trained on Chrome data and tested across all browsers. The results demonstrated high accuracy, F1-score, precision, recall, and minimal log-loss error for Chrome and Firefox, but subpar performance on Brave, potentially due to its distinct data distribution and feature set. The research suggests that these classifiers are viable for detecting web trackers in Chrome and Firefox. However, real-world application testing remains pending, and the distinction between tracker types and broader label sources could be explored in future studies.
Web Privacy, Web Tracking, Web Privacy Measurement, HTTP/S Response Header, Machine Learning
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
The rapid evolution of the Web has cemented its role as an integral component of modern life. Through desktop and mobile browsers, users access a vast array of information and services. However, every interaction, from content consumption to mere site visitation, reveals information about the habits and preferences of users, which in turn has implications for personal privacy and online security.
These privacy concerns are highlighted by the ubiquity of behavioural targeting. By analyzing a user’s browsing history, a detailed profile can be constructed, facilitating the delivery of tailored advertisements and content. This profiling is made possible by web tracking that employs sophisticated techniques to track users across the Web. Given the increasing concerns about such practices (Melicher et al., 2016; Chanchary and Chiasson, 2015), there is a growing emphasis on research in web privacy.
An established countermeasure against unsolicited HTTP communications is the use of filter lists (Traverso et al., 2017), such as EasyList (eas, 2023a) and EasyPrivacy (eas, 2023b). However, their dependency on manual curation makes them vulnerable to oversights and rapid obsolescence (Hashmi et al., 2019; Alrizah et al., 2019; Fouad et al., 2018). Moreover, some trackers employ strategies like domain masking to evade these lists (Chen et al., 2021).
In an attempt to mitigate these shortcomings, recent efforts have explored the potential of Machine Learning (ML) in this domain. Various classifiers, some achieving accuracy metrics in the 80th percentile (Cozza et al., 2020; Iqbal et al., 2021), have been developed using diverse data sources such as HTTP headers and JavaScript calls. Yet, a majority of these focused primarily on HTTP requests, potentially neglecting other informative data streams. Additionally, the predominant reliance on Firefox for data collection may limit the comprehensiveness of these studies.
Notably, there seems to be a gap in literature regarding the use of HTTP response headers as the sole data source for tracker classification across diverse browsers, despite studies in the related field of online security highlighting the efficacy of this approach (McGahagan et al., 2019). This study intends to bridge this gap. We strive to develop classifiers adept at web tracker detection, while confronting challenges such as data sparsity, high dimensionality, and class imbalance.
Our key contributions are:
  1. (1)
    We conduct a rigorous study of HTTP response headers across multiple browsers, focusing on their potential for web tracker detection within the Tranco top 10k websites. Furthermore, we evaluate dataset similarities to discern differences in HTTP response headers across the assessed browsers.
  2. (2)
    We present a semi-automated ML pipeline which facilitates the training of eleven classifiers tailed for web tracker detection. Our findings indicate that selected tree and gradient boosting models, when applied to a subset of binary-encoded HTTP/S headers, are capable of achieving ROC-AUC, AUPCR, and F1 scores exceeding $0.9$.
  3. (3)
    We conduct a comprehensive and multi-faceted evaluation of these classifiers, which includes longitudinal studies to gauge potential concept-drift and evaluates the cross-browser performance of classifiers trained solely on traffic data collected within the Chrome browser. Notably, our evaluation incorporates thirteen metrics, offering a holistic view of the binary classification challenge.
  4. (4)
    We assess calibration effects on final classifier performance — a consideration seldom addressed in prior research. This allows us to discern if models trained solely on data from one browser can be effectively deployed on data from alternate browsers without the necessity for retraining.
The remainder of this paper is structured as follows. Section 2 reviews past research in the field of web tracker detection. Section 3 introduced definitions, our research questions, and the analyzed data sets. Section 4 explores HTTP/S response headers from Chrome, Firefox, and Brave data sets, providing general findings and identifies potential challenges for the ML pipeline. Section 5 describes the design and implementation of the ML pipeline, including selected models, metrics, and baselines. Section 6 presents the evaluation of the trained classifiers. The penultimate Section 7 discusses our results and limitations. Section 8 concludes the paper.
2. Related Work
This section provides an overview of existing research on web tracker detection based on machine learning. To highlight how our proposed classifier differs from prior work, we focus this overview on classifiers that rely solely on HTTP protocol information. For the sake of completeness, however, we summarize other approaches at the end of the section.
2.1. Classification Based on HTTP Information
Approaches to the detection web trackers with the support of machine learning date back to the early 2000s. In 2010, Yamada et al. (Yamada et al., 2010) were the first to use temporal link analysis to build a graph based on hosts related to HTTP requests from the traffic generated by corporate network. Their classifier considers temporal information such as the visiting time of a specific website or host. Yamada et al. consider this a discriminating criterion based on the assumption that users spend less time on websites of web trackers, as trackers generally appear as third parties that are contacted briefly in the context of other websites rather than first parties that users actively visit.
In 2013, Bau et al. (Bau et al., 2013) described and discussed the usefulness of machine learning as an effective counter to web tracking. They outlined the problems with manually curated filter lists - problems that still hold true today, such as the reliance on manual curation and the inflexibility of static lists. Beyond building an Elastic Net model focused on traces of content extraction in the Document Object Model and HTTP headers to show the advantages of automated approaches, Bau et al. laid down requirements for future machine learning models as well as problems facing researchers in this domain.
A year later, Bhagavatula et al. (Bhagavatula et al., 2014) tried to identify ad-related URLs, a topic closely connected to the web tracking detection problem. To this end, they crawled the Alexa top 500 websites to compile an extensive dataset with six distinct feature sets based on the URLs themselves as well as additional metadata. Their analysis approach is based predominantly on Support Vector Machines (SVMs) with different kernels, achieving average accuracy and precision scores above 89%. Like Bau et al. before them, Bhagavatula et al. recognized the need for machine learning models to replace filter lists and put forward additional requirements for production-ready machine learning approaches, such as optimized training times.
In 2015, Gugelmann et al. (Gugelmann et al., 2015) proposed an alternative approach by deriving several statistical features from HTTP traffic, e.g. up- and download volume. In contrast to prior studies, they thoroughly described their model-building process and emphasized their feature engineering. Their feature selection process included the use of Pearson’s correlation coefficient as well as the consideration of a feature’s entropy. In addition, they employed an unspecified classification tree, among other classifiers, to find an optimal Naïve Bayes model that achieved a precision of 83% and was able to correctly identify previously unseen trackers.
In the same year, Li et al. (Li et al., 2015) proposed TrackAdvisor, a system to classify HTTP requests based on the collective statistics of cookies contained therein. Using FourthParty, they captured 563,031 requests and identified 22,270 cookies that were sent to third parties. Rather than relying solely on filter lists, they used a multi-step process with an emphasis on manual inspection to label the captured data. Notably, the training and test sets created from the captured data comprised only 500 requests each and had a balanced class distribution, which contrasts with the distribution found in the original data. Nevertheless, Li et al. used these sets to train a SVM that ultimately achieved a precision of 99.4% and a recall of 100%. Like Gugelmann et al., they point out the usefulness of the HTTP Referrer header for third-party tracker detection.
In 2016, Dudykevych and Nechypor (Dudykevych and Nechypor, 2016) also used cookies as the basis for their classifiers. In contrast to Li et al., however, they developed their own crawler, had a larger data set of 550 000 cookies, and developed a different feature set. The latter is only described briefly, but information from the cookies themselves and the HTTP requests were used as features. It can be assumed that the features are numerical and categorical, as the authors used Principal Component Analysis and correlation, which require numerical data, as well as binary encoding of categorical features. The resulting classification accuracies ranged from 82.01% for Naïve Bayes to 95.43% for the regularized logistic regression method.
The analysis of HTTP traffic for and with machine learning models is also an important topic in web security research, albeit with a different target (i.e. label) than web privacy research. In a 2019 study, McGahagan et al. (McGahagan et al., 2019) analyzed response headers from 45 874 websites and extracted 672 features, resulting in a final feature vector with 22 properties. In addition, they tested whether over- and undersampling had any effects on the model’s performance. The coverage of metrics, models, different training and tuning combinations, and extensive feature engineering documentation can be considered standard among machine-learning practitioners but is somewhat rare among prior related research. An important contribution of this paper is the in-depth presentation of the usefulness of HTTP response headers for automated ML models.
Unlike most papers discussed in this section, which have used data from the Firefox browser to test the performance of supervised learning models, Din et al. (Din et al., 2020) used Deep learning (DL) techniques for Brave and other Chromium-based browsers. Although their approach is geared towards the image-based detection of advertising in websites rather than web tracking, their incorporation of multiple browsers is highly relevant for models that classify web trackers.
Cozza et al. (Cozza et al., 2020) developed a system that not only classifies HTTP requests but also automatically updates the filter list. Although this study is not the first to detect previously undocumented trackers using machine learning techniques, the deployment of a tracker-detecting browser extension that combines a filter list with a machine learning models and updates its ground truth marks an important step forward in tracker detection technology.
Guarino et al. (Guarino. et al., 2020) built upon prior research by Gugelmann et al. and Li et al. by training multiple model based on combinations of features used by these studies, including different HTTP headers and the ratio of upstream and downstream data transmissions. Their top performer is a Random Forest model with an average accuracy of 92%. It should be noted, however, that this study is based on a relatively small dataset of just 1000 observations. Nevertheless, the authors astutely emphasize that content delivery networks (CDNs) are a primary cause of false positives in the domain of web tracker detection due to similarities in traffic patterns between CDNs and third-party trackers.
Taking advantage of the observation that HTTP-based communication between clients and servers often comprises cascades of multiple HTTP messages, Iqbal et al. (Iqbal et al., 2022) conducted an extensive study that uses entire request chains rather than individual requests as the basis for classifying web tracking. Notably, they were among the first to consider responses as well as requests by developing three distinct feature sets: the first is based on the entire request chain, the second relies on information pertaining to individual requests and the third considers response headers and metadata. By training a Random Forest model on five distinct datasets with different configurations from separate crawls, Iqbal et al. achieved accuracies above 98%. However, if the model was trained on only one crawl and tested on the others, the accuracy dropped by up to 14%, highlighting the difficulties of applying models to datasets with varying configurations, e.g. data collected across different browsers.
2.2. Other Detection Approaches
Beyond the HTTP-focused approaches outlined above, a variety of classification methods utilizing different data sources for the extraction of distinctive criteria have been proposed and evaluated. These include, classifiers focused on JavaScript code and API access events that aim to identify device fingerprinting techniques, pure URL-based classifiers, CNAME cloaking detectors, and graph-based classifiers that leverage the topology of the Web to identify third-party tracking in particular. In the following paragraphs, we will briefly highlight notable research on these alternative approaches.
Focusing on JavaScript API calls, Wu et al. (Wu et al., 2016) developed DMTTrackerDetector, a system using decision trees in which observed API calls serve as features. The system, comprising a 717-dimensionial feature vector, achieved 97.8% accuracy and addressed obfuscation issues, showing the distinct JavaScript API usage by trackers and non-trackers.
Bahrami et al. (Bahrami et al., 2021) refined the scope of their analysis by focusing on device fingerprinting rather than API calls in general. Their approach combines graph analysis with supervised and unsupervised learning to detect web trackers in historical JavaScript file data collected between 2010 and 2019 by creating a graph that includes temporal aspects and the output of AST-based keyword extraction. Their method, which includes nine graph-related features and graph embedding, as well as a clustering model, yielded accuracies up to 89.13% and identified distinct clusters for fingerprinting and functional APIs, highlighting the viability of such an approach.
Kalavri et al. (Kalavri et al., 2016) also proposed a graph-based approach, using neighborhood analysis and label propagation on a bipartite graph representing connections between first-party websites and third-party services to detect web trackers, achieving over 97% accuracy. They analyzed six months of data, finding that most trackers are closely linked, with precision scores ranging from 64 to 92%.
Combining these graph-based approaches, Castell-Uroz et al. (Castell-Uroz et al., 2021) proposed a tripartite network graph to identify third-party tracking resources and their origin domains within first-party websites. By applying this approach to a dataset comprising over 90 thousand domains, they were able to detected new trackers through their hash value popularity and dirt level, finding that higher values correlated with the likelihood of a given resource to be tracking-related.
Conversely, Metwalley et al. (Metwalley et al., 2015) developed an unsupervised algorithm analyzing URL queries and HTTP request headers for identifying trackers. Their method, applied to 200 Alexa-ranked websites, identified 34 new trackers and successfully supported the analysis of cookie-matching practices, though it had limitations in detecting all known trackers.
Finally, Yu et al. (Yu et al., 2016) proposed a browser extension using k-Anonymity principles to classify third-party trackers based on data sent to users. Leveraging data from 200,000 German users of the Cliqz browser, their system outperformed traditional blockers like Disconnect, which had a 38.8% false positive rate.
2.3. Summary
Despite over a decade of dedicated research, web tracking is as prevalent as ever, new tracking methods are developed, and the most widespread form of protection, namely blockers and their filter lists, exhibit several key problems. Filter lists, which are also used as the ground truth for labeling in many studies related to web privacy, are often maintained by a small group of users, contain irrelevant rules, are rather slow to add new rules, and can simply be circumvented by changing URLs. In response, researchers have developed a wealth of automated approaches, utilizing HTTP messages, JavaScript, or the HTML files of websites themselves to classify third parties as trackers.
A crucial component of this ongoing research effort is the availability and quality of data on which to train and evaluate machine learning approaches.
Despite this, HTTP response headers are often neglected in this regard. There are cases where authors document the collection of HTTP responses (Acar et al., 2013) but not their actual use. In other cases, the data is collected, but only used in support of other approaches such as the classification of cookies (Lerner et al., 2016; Englehardt et al., 2015).
Additionally, the majority of studies are based on datasets collected using just one browser, most commonly Firefox. Cassel et al. (Cassel et al., 2022) discuss the need for cross-browser examinations to avoid biased datasets. This is supported by research that shows that characteristics such as JavaScript API calls differ between browsers (Yang and Yue, 2020).
In summary, our review of related research has shown that, although the application of machine learning to the field of web tracking identification and classification has been investigated in numerous studies, a research gap still exists with regard to cross-browser performance of classifiers and the usefulness of HTTP response data for classification purposes.
3. Approach
In this study, we address the research gap identified in our review of related research by evaluating the (cross-browser) performance of commonly used machine learning classifiers that are trained on HTTP/S response headers. In this section, we discuss our underlying research questions and present the approach we used to address them.
3.1. Preliminaries
A core principle of automated web tracker detection is the definition of what constitutes a tracker. The focus hereby lies on the tracking entities and not on the activity of tracking. Our definition of what constitutes a tracker or non-tracker is coupled to our ground truth to label each dataset. We define them as follows:
[…]
$\text{match}(h,r)$ :$:\begin{cases}1&\text{if }h\text{ matches rule }r\\ 0&\text{otherwise}\end{cases}$
$T$ =$=\{h\in H\mid\exists r\in R:\text{match}(h,r)=1\}$
$NT$ =$=\{h\in H\mid\forall r\in R:\text{match}(h,r)=0\}$
In other words: if the remote hostname (i.e. the sender) of a response matches at least one rule in our aggregate list compiled from the EasyList and EasyPrivacy filter lists, we label the response as a tracker.
For the sake of simplicity, we will use the terms header and response header interchangeably to refer to HTTP/S response headers throughout the remainder of the paper, unless explicitly stated otherwise.
3.2. Research Questions
The main goal of this study is to analyze whether response headers are useful for automated web tracker detection. Based on the ML approaches of prior research and the observed areas to further extend our empirical understanding and detection of web trackers, we formulate the following research questions that guide us throughout this study:
  1. (1)
    What are the differences and similarities between web trackers and non-trackers with regard to their response headers?
  2. (2)
    Can trained classifiers detect web trackers with only response header features while achieving an F1-Score higher than 0.9 in an imbalanced setting?
  3. (3)
    How large is a classifier’s performance degradation when it is trained on one browser’s HTTP traffic data and then applied to other browsers?
To answer these research questions, we develop a semi-automated ML pipeline that transforms HTTP traffic datasets collected with T.EX (Raschke et al., 2019) and uses it to train several supervised binary classifiers. We use reliable data that was compiled by prior research to facilitate future comparisons between new classifiers.
To test the robustness and reliability of our classifiers, we train them on HTTP data from Chrome and apply them to different datasets with a variety of configurations, including other browsers (latitudinal comparison) as well as different time frames within the same browser (longitudinal comparison). Furthermore, past research observed an uneven distribution of trackers and non-trackers, which is why we train our classifiers with an imbalanced distribution, to assess how effective they perform with the likely true distribution.
$\text{Chrome}_{22}$ $\text{Firefox}_{22}$ $\text{Brave}_{22}$ $\text{Chrome}_{23}$
T NT T NT T NT T NT
#Responses 256 202 602 516 309 391 526 865 2969 561 694 206 747 596 066
#Responses in % $\approx 0.3$ $\approx 0.7$ $\approx 0.37$ $\approx 0.63$ $\approx 0.5$ $\approx 99.5$ $\approx 0.26$ $\approx 0.74$
#Unique headers in total 2213 7577 2516 7322 237 7348 2164 7398
Unique Headers per response ($Q1,\tilde{\mu},Q3$) 10,13,16 12,15,19 12,15,18 13,16,19 13,18,20 12,15,19 10,13,16 13,15,19
Table 1. Overview of the collected data sets from (zen, 2023), split by data sets as well as tracker (T) and non-trackers (NT).
Figure 1. This figure illustrates the usage headers across four different browsers. The left plot (a) displays the frequency distribution of headers for each browser, showing how often each header is present in the dataset. The right plot (b) presents a boxen plot of the same data, which provides a more detailed view of the distribution, revealing variations and trends among the browsers.
3.3. Data Collection
We used existing data sets, containing labeled HTTP response headers by EasyList and EasyPrivacy. Following suggested good practices by (Demir et al., 2022), the crawl should fulfill three citeria: (i) repeated crawls across multiple browsers, namely Chrome, Firefox, and Brave, (ii) the documentation of the geographical region as well as the technical environment, and (iii) web sites from the Tranco ranking as it proved to be less biased than the Alexa ranking (Pochat et al., 2018). A previously crawled data set by Raschke using T.EX - The Transparency EXtension, fulfills all three criteria and is described as follows:
Multiple simultaneous and stateful crawls (6 x Chrome, 6 x Brave, 6 x Firefox) of the Tranco top 10K websites (as of 16th of May 2022) were performed on the 12th of August 2022 with T.EX […]. Measurements had been carried out on 18 Amazon Web Services instances (c5.large) running Windows Server. All instances were launched in Frankfurt, Germany (eu-central-1).
After completing the crawls, each instance’s extension storage was extracted […]. Note: 2 of the 6 Firefox instances crashed during the crawl. Therefore, only 4 Firefox datasets are available (zen, 2023).
To assess a potential concept drift and therefore model degradation over time, we crawled one more data set from 2023 for Chrome. We followed a similar methodology as described in (zen, 2023), using T.EX with Chrome to explore the top 10K websites from the Tranco ranking as of March 29th, 2023.
Our reasons to choose only Chrome for a new data collection are as follows: Chrome is the most widely used in the real world with a market share of 66.13% on desktops, whereas Firefox has 7.1% (sta, 2023) and Brave has less than 1% (kin, 2023). Based on this distribution, the Chrome data set serves as the foundation for training to reflect and reach the majority, i.e. using data that most users encounter themselves. In addition, web privacy research often relies on a simulated Firefox browser with Selenium due to the use of OpenWPM, leaving space for other browsers of interest. In addition, it helps us to answer the question whether a model based on Chrome data can be applied to Firefox or Brave while achieving high performance metrics?.
4. Initial Response Header Analysis
To get a better understanding of the data underlying our analysis before designing our machine learning pipeline, we conduct an initial analysis of the data, the results of which we present in this section. We first explore response headers across all three browsers to find empirical evidence of their potential relation to trackers. Furthermore, we identify the characteristics of these datasets to apply the appropriate methods in our ML pipeline. We use these insights to develop our classifiers to evaluate the effectiveness of response headers for web tracker detection.
4.1. Descriptive Analysis of Headers
A descriptive analysis is useful to gain an understanding of the data structure and the distribution of classes and features within the dataset. Table 1 summarizes descriptive metrics for each dataset. We start with the datasets $\text{Chrome}_{22}$, $\text{Chrome}_{23}$, and $\text{Firefox}_{22}$. All three datasets show similar results for the distribution between trackers and non-trackers, the number of unique headers in total and per response. For example, we can observe a small quantitative change in the number of unique headers between $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$. $\text{Chrome}_{22}$ lists a total of 7577 unique headers, while $\text{Chrome}_{23}$ has a slightly lower total of 7398 for non-trackers. When comparing trackers, $\text{Chrome}_{22}$ had 2213 unique headers, which is nearly the same as $\text{Chrome}_{23}$’s count of 2164. On average, non-trackers have more unique headers as well as a higher median $\tilde{\mu}$ and interquartile range than trackers.
The $\text{Brave}_{22}$ browser with its standard configuration is the only one that has a low number of trackers and unique headers in tracking-related response messages. Interestingly, the $\tilde{\mu}$ equals 18 for trackers and 15 for non-trackers, which represents the only instance where trackers had a higher number. Although the dataset has almost only non-trackers, the number of unique headers is similar to the other datasets.
Figure 1 shows a frequency plot, which is a distribution graph that shows how often each header appears in the dataset. It uses a logarithmic scale to manage the wide range of frequencies, which is a common approach when data spans several orders of magnitude. This plot typically demonstrates a power-law distribution, characterized by a steep drop-off where a small number of headers are occurring frequently, while the majority are rare. This type of distribution is common in HTTP data, where few headers dominate, and many others have a long tail of low frequency. This sparseness indicates that while there are many different headers, only a few are consistently used across different websites and by different browsers.
The boxen plot graph of Figure 1 offers a visual representation of the distribution of header frequencies across different browsers and versions. The spread and median values suggest that the distribution of headers is not uniform across browsers, with some browsers having a higher tendency to use certain headers than others. Furthermore, if the median line in the boxen plot is closer to the bottom of the box, it suggests a right-skewed distribution, aligning with what the frequency plot shows — a few headers with high frequency and many with low frequency.
Figure 2. The Venn diagram reveals that a core set of 5672 headers are common across Chrome, Firefox, and Brave browsers across two years. Unique headers are present in each browser, with $\text{Chrome}_{23}$ having the highest number of unique headers, which may reflect browser-specific features or experimental headers. Firefox had a higher header similarity with Chrome than with Brave.
Figure 2 represents a Venn diagram and provides a direct comparison of response headers between different browser versions, highlighting data and feature similarity. The intersection numbers, such as the 5672 headers common to all browsers, suggest a significant overlap in features across different platforms. However, the presence of headers unique to each browser or shared by only two browsers (e.g., 1427 unique to $\text{Chrome}_{22}$ and 154 unique to $\text{Brave}_{22}$) indicates that while there is a core set of features, there is also a notable amount of variation which can be critical for web tracker detection.
The Empirical Cumulative Distribution Function plot for $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$ in Figure 3 row (A) show that the Content-Length values for non-trackers are generally higher than those for trackers, which suggests that non-trackers tend to send larger responses. This is consistent across both browser versions, indicating that while browsers evolve, certain characteristics like the smaller size of tracker responses remain consistent. The second row (B) plots the feature similarity for the X-XSS-Protection header, showing that header values can be very similar between the datasets, i.e., the training and test sets. It further supports our observation that the datasets are similar and that trackers and non-trackers have differences across the usage of HTTP header fields and values. Although we present only few selected examples here, they should highlight the general challenges that come with response messages datasets.
Figure 3. Row (A) presents the ECDFs for both Chrome datasets for the Content-Length header values. We set a cut-off value at 10 000 to highlight the core observation and more than half of the responses have a value are below this threshold – $\tilde{\mu}_{T22}=62$, $\tilde{\mu}_{NT22}=8068$ and $\tilde{\mu}_{T23}=45$, $\tilde{\mu}_{NT23}=8021.5$. Row (B) shows how similar the values across Chrome, Firefox, and Brave are for the X-XSS-Protection header.
Additional characteristics of the datasets include the high amount of missing values per header. One explanation might be that servers do not need to use every header field that exists, and it is common practice to use developer headers that are not part of the standard, i.e. they are not registered in the Request for Comments 4229111RFC 4229, https://www.ietf.org/rfc/rfc4229.txt, accessed: November 30, 2023. There are multiple headers that contain mostly unique values, e.g., Date. Lastly, according to the used filter lists, most HTTP response messages are not related to any tracking activities. This entails a class imbalance to be handled carefully to derive meaningful results from the final models.
5. Methodology
Based on our findings from the data exploration in Section 4, we propose our methodology to assess the effectiveness of classifiers based on response headers.
5.1. Problem Formulation
We consider the problem of binary classification, where we aim to predict the presence of web trackers based on response headers. The headers, represented by the feature vector $x\in\mathbb{R}^{d}$, are nominal and can be both discrete and continuous variables. These variables are derived from HTTP headers of four different browser types, denoted as the set $B=\{\text{Chrome},\text{Firefox},\text{Brave}\}$. For this study, we specifically consider versions $\text{Chrome}_{22},\text{Chrome}_{23},\allowbreak\text{Firefox}_{22},\text{Brave}_{22}$.
Given a training set $\mathcal{D}_{\text{train}}=\{(x_{1}y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}$ is a feature vector of the $i$-th HTTP header and $y_{i}\in\{0,1\}$ represents the associated class label (with ’1’ indicating the presence of a web tracker), our goal is to train a set of classifiers $f_{j}:\mathbb{R}^{d}\rightarrow\{0,1\}$ for $j\in M$ that minimizes the empirical risk:
(1) $R(f_{i})=\frac{1}{N}\sum_{i=1}^{N}L(y_{i},f_{i}(x_{i})),\quad j\in\{1,\ldots,M\},$
where $L(y_{i},f_{i}(x_{i}))$ is the loss incurred by predicting $f_{i}(x_{i})$ when the true label is $y_{i}$. The specific form of the loss function $L(\cdot)$ will be dependent on the particular classifier $f_{j}$ used.
We consider datasets represented by $D_{s}$, where $s$ can pertain to training, testing, validation, or calibration subsets of the data. We observe several challenges from our data exploration:
  1. (1)
    High Cardinality: The feature vectors $x_{i}$ in our datasets $D_{s}$ can exhibit high cardinality for certain features. This problem can pose challenges in terms of model training and encoding.
  2. (2)
    Sparseness: The feature vectors $x_{i}$ in our datasets $D_{s}$ are represented as sparse vectors, with each vector defined as:
    $x_{i}=\{v_{1},v_{2},\ldots,v_{d}\}$
    Here, a significant proportion of the entries $v_{j}$ within $x_{i}$ are either zero or represent missing data, leading to the overall sparseness of $D_{s}$. Such sparseness can introduce computational complexities and may challenge the performance of certain machine learning algorithms not optimized for sparse data.
  3. (3)
    Class Imbalance: For datasets collected from a given browser $B$ and represented as $D_{s}$, the class distribution can exhibit imbalance. The majority class is denoted by $\alpha_{B}$, and the minority class is represented by $\beta_{B}$, satisfying $\alpha_{B}+\beta_{B}=1$. The degree of this imbalance can vary based on the browser as shown in Table 1, with some showing a more pronounced disparity between $\alpha_{B}\text{ and }\beta_{B}$. Such imbalance can introduce bias to models, favoring the majority class and potentially undermining the accurate prediction of the minority class.
  4. (4)
    High Dimensionality: With a high dimensionality $d$ and a high number of observations $n$ per dataset, the classification is susceptible to the Curse of Dimensionality, potentially leading to overfitting and necessitating intricate dimensionality reduction techniques. We have to apply a dimensionality reduction process that achieves a reduced feature space $X_{i}^{\prime}\subseteq X$, entailing a reasonable training time.
  5. (5)
    Data Similarity: The datasets $\mathcal{D}_{s}$ collected across different browsers can exhibit pronounced similarities in their feature distributions and values. Given the high dimensionality of our feature space, exhaustive pairwise comparisons can be computationally expensive and time-consuming, highlighting the challenge of accurately gauging data similarities in such a complex space.
5.2. Machine Learning Pipeline
Our machine learning pipeline, which we present in the following, addresses the previously stated problems for the classification of web trackers. Many different approaches have been applied in order to uncover web trackers. Another closely related field to web privacy is web security research. Although they seek to detect different entities, both have the exploitation of HTTP data in common. Typically, a dataset goes through extensive data processing and feature engineering to find useful features, which includes a thorough analysis of the values within each feature. Despite this common conception, Laughter et al. (Laughter et al., 2021) were interested if features that merely represented the presence of an HTTP header in a binary fashion were usable in a classifier to detect malicious HTTP requests based on HTTP headers and URLs. In the end, some of their classifiers performed well with accuracies above 90%. Similar results were observed in (Bhagavatula et al., 2014) and partly by (Dudykevych and Nechypor, 2016).
This method is interesting on two levels: (i) it cuts down the time of feature engineering, which is fairly complex in this use case as HTTP headers are mainly categorical data, and (ii) it should help when the models are deployed and used in production, e.g. if HTTP data is received, it should be evaluated in real time and checking the presence of headers is faster than elaborate processing before evaluation. Therefore, the binary representation of HTTP headers as a feature will be tested.
First and foremost, we train our classifiers based on the $\text{Chrome}_{22}$ dataset and use all other datasets as test sets, which is in accordance with RQ3. Generally, each step in our pipeline focuses on reducing the dimensions of the datasets, as most headers are rarely used, thus reducing potential overfitting. This approach further reduces the sparseness and potential features with high cardinality. The latter challenge is solved by employing the binary-encoding method as it flattens the respective feature to two possible values.
The sparseness and high dimensionality are related to each other, in a sense that most headers and their corresponding values are rare or even unique. We apply multiple methods to reduce the dimensionality: (i) a low variance filter to remove headers with only one value, (ii) the removal of headers with a missing value ratio of one, (iii) headers related to only one label are dropped, and (iv) Approximate string matching, sometimes referred to as Fuzzy Matching to combine header fields with a similarity of at least 80% based on a weighted ratio of multiple distance metrics, e.g., Damerau Levenshtein or Hamming. All of this reduces the sparseness further and binary-encoding simplifies the sparse matrix.
Although a multitude of different resampling techniques exists to combat class imbalance, the question of whether this step is actually necessary remains. Some disadvantages of such methods are, for example, the removal of observations, the replication of existing observations, or the creation of synthetic samples. Selecting and applying a sampling method can lead to overfitting and overoptimism as well (Santos et al., 2018), notwithstanding the question of whether it is even feasible to train the models on a distribution that was not observed in real-world data. Our pipeline preserves the observed distribution as the imbalance is mild (goo, 2023) in the $\text{Chrome}_{\textbf{22}}$ dataset, i.e., a sufficient number of samples for each class exists, which allows proper training to differentiate between classes. Furthermore, we compute metrics that consider or are not affected by the imbalance of the datasets. The same applies for most of the selected ensemble models.
Fuzzy matching is especially useful due to the presence of hundreds of similar headers within our data. During our initial data analysis deiscussed in Section 4, we identified many examples where headers with similar purpose were written or formatted differently, e.g. content-lenght, cteonnt-length, or ntcoent-length, instead of content-length. Accordingly, simply comparing the header names is not enough, which is why we automatically assess the value similarity to reduce potential mismatches.
Lastly, we apply Isotonic Regression instead of Platt Scaling to calibrate the probabilities outputted by our models, following the results from (Niculescu-Mizil and Caruana, 2005). This is done by fitting an Isotonic Regression model to the models’ predictions and the true labels, and using this to adjust the predicted probabilities. The calibration function for model $i$ is denoted as $C_{i}(f_{i}(x))$.
We summarise our pipeline in Figure 4, including the data collection. Note that we did not apply any data imputation to keep the final datasets as close to the original collected datasets as possible. A validation set for hyperparameter tuning was not created because we want to assess the effect of calibration on its own. Moreover, tuning requires substantially more computational power, e.g. by using Grid Search.
Figure 4. The parameterized pipeline automatically processes the collected datasets and trains the classifiers.
5.3. Classification Models
As discussed in Section 2.1, prior studies (Cozza et al., 2020; Li et al., 2015; Bhagavatula et al., 2014; Gugelmann et al., 2015) often used similar classifiers despite the broad range of available models. We build upon these established selections, refining our choice of models based on three conditions: (i) the models should be trainable in a reasonable amount of time, (ii) they should be applicable to this specific use case, and (iii) they should preferably be ensemble methods, although we include a handful of non-ensemble models that were used in past research to allow for direct comparisons between our results and those of related studies. We prefer ensemble methods as they usually perform better than their counterparts (Dietterich, 2000), leading to their popularity over the past few years. Nevertheless, to our best knowledge, they have not been widely used in past web privacy research, except for Random Forests (RF). In Table 3, we provide an overview of our selected models and their use in related work.
Model Ensemble Past Research
Decision Tree No (Gugelmann et al., 2015; Dudykevych and Nechypor, 2016; Raschke et al., 2023)
Random Forest Yes (Cozza et al., 2020; Wu et al., 2016; Guarino. et al., 2020; Iqbal et al., 2022; Raschke et al., 2023), (Dudykevych and Nechypor, 2016) (U), (McGahagan et al., 2019)
Extra Trees Classifier No Not used
Logistic Regression No (Bhagavatula et al., 2014; Gugelmann et al., 2015; Dudykevych and Nechypor, 2016; Wu et al., 2016; Raschke et al., 2023), (McGahagan et al., 2019)
Naïve Bayes (Gaussian or Bernoulli) No (Yamada et al., 2010; Bhagavatula et al., 2014; Gugelmann et al., 2015; Dudykevych and Nechypor, 2016; Wu et al., 2016; Guarino. et al., 2020)
Gradient Boosting Yes (McGahagan et al., 2019), (Raschke et al., 2023)
Light-Gradient Boosting Machine Yes Not used
Adaptive Boosting Yes (Raschke et al., 2023)
Hist-Gradient Boosting Yes Not used
eXtreme Gradient Boosting Yes (Raschke et al., 2023)
Table 2. Overview of the selected machine learning models and the number of times they were used in past research from Section 2.1. One paper did not exactly specified which ensemble model was used, thus, the most likely model was marked with Unspecified (U). Papers from web security are underlined.
Other supervised learning methods, e.g., Multi-layer Perceptrons (Cozza et al., 2020; Guarino. et al., 2020), KNN, and especially SVM, could have been used. The latter was successfully applied in multiple papers (Cozza et al., 2020; Li et al., 2015; Bhagavatula et al., 2014; Guarino. et al., 2020). However, both suffer in terms of time and computational complexity, e.g. SVM has a train time complexity of $O(n^{2})$ where $n$ is the number of observations (Steinwart and Christmann, 2008). In addition, deep-learning models are not considered as the focus is on supervised classifiers and simple rather than complex models. A graph-based approach, used by several papers (Kalavri et al., 2016; Castell-Uroz et al., 2021) will not be pursued because it does not fit well with the targeted approach, and direct connections between individual HTTP responses will not be analyzed, e.g. through grouping based on their URL or building a graph-based on their order.
5.4. Baseline Models
We compare our classifiers against three baseline models and multiple metrics that we define in Section 5.5. We select one tree classifier, the Decision Tree (DT), that is also the building block of six selected classification models. Two simplistic and relatively robust baselines, namely Naïve Bayes (NB) and Logistic Regression (LR).
Additionally, it is good practice to compare against a baseline approach that differs from the presented classifiers while using the same dataset. Raschke et al. (Raschke et al., 2023) presented a graph-based approach named T.EX-Graph where edge and node attributes as well as centrality metrics served as features. Furthermore, the authors included information from the responses but did not pursue classifiers that are solely based on response headers. This comparability should help in assessing the performance gains and the relevance of response header values for automated web tracker detection. However, the authors applied Synthetic Minority Oversampling to balance their dataset. We focus on their results using the FQDN dataset, where hosts are modeled as fully qualified domain names (FQDN) because the tracker distribution is similar to ours with a NT to T ratio of 68:32.
5.5. Evaluation Metrics
We measure the performance of our classifiers according to nine different metrics. Several metrics are common in the web tracking literature, e.g., accuracy, precision, or F1-Score, while others were selected due to their applicability to our classification problem and to understand the relationship with the underlying data in more detail. Additional metrics are reported in the Appendix, e.g., the confusion matrix because its interpretation is already captured by other metrics more concisely. Moreover, we calculate confidence intervals (CI) for these metrics using bootstrap sampling with 599 samples (Wilcox, 2010).
Metric Past Research
(Avg.) Accuracy (Gugelmann et al., 2015; Dudykevych and Nechypor, 2016; Raschke et al., 2023; Cozza et al., 2020; Guarino. et al., 2020; Iqbal et al., 2022; Bhagavatula et al., 2014), (McGahagan et al., 2019)
Precision (Gugelmann et al., 2015; Raschke et al., 2023; Iqbal et al., 2022; Bhagavatula et al., 2014; Li et al., 2015)
Recall (Gugelmann et al., 2015; Raschke et al., 2023; Iqbal et al., 2022; Li et al., 2015)
F1-Score (Gugelmann et al., 2015), (Raschke et al., 2023)
AUC-ROC (Gugelmann et al., 2015), (Bhagavatula et al., 2014), (McGahagan et al., 2019)
Confusion Matrix (Guarino. et al., 2020)
MCC (McGahagan et al., 2019)
AUCPRC Not used
Log-Loss Not used
Table 3. Overview of performance metrics and the number of times they were used in past research from Section 2.1.
The first metrics that are commonly used in web privacy research are Accuracy, Precision, Recall, False Positives (FP), True Positives (TP), False Negatives (FN), and True Negatives (TN). Two accuracy scores can be calculated, one from the perspective of trackers, the positive label, and one from non-trackers, the negative label. However, accuracy is not particularly useful with imbalanced datasets because a model could always predict the majority class and achieve accuracy with the same value as the distribution of said class. We calculated additional, more reliable metrics to quantify the performance of a classifier with an imbalanced dataset as follows:
Balanced Accuracy:
Broderson et al. (Brodersen et al., 2010) introduced the Balanced Accuracy (BACC) metric for imbalanced datasets in 2010 as an alternative to average accuracy scores, which are, according to the authors, problematic for confidence intervals (CI) and imbalanced datasets. The metric comprises recall and TNR and is defined as:
(2) $BalancedAccuracy=\frac{1}{2}\left(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\right)$
F1-Score:
Another metric that is commonly used for classification models is the F1-Score which expresses a model’s performance in one value, despite the ambivalent nature of precision and recall. Compared with BACC, it emphasizes positives more, whereas BACC takes negatives more into account, which can result in two very different results. The F1-Score is mathematically defined as the harmonic mean between recall and precision:
(3) $F_{1}\text{-}Score=\frac{2\cdot Precision\cdot Recall}{Precision+Recall}$
Log-loss Score:
This metric, sometimes referred to as cross-entropy-loss, is particularly useful for classification problems as it focuses on the predictive performance of a model. The function is applied to each observation individually, and then the average over all elements is calculated. It describes how large the error between a prediction probability and the true label is, and a value closer to zero represents a very good score. However, this metric can be influenced by the imbalance in a dataset, i.e. the baseline log-loss gets smaller with an increase in imbalance, and the log-loss metric should always be compared to its baseline value, e.g. the baseline log-loss for a dataset with a 70:30 ratio equals 0.611, and 0.325 for datasets with a 90:10 ratio (tds, 2023). All in all, the model’s log-loss should be lower than its baseline. The math formula is defined as:
(4) $L_{log}(y,p)=-[y\cdot log(p)+(1-y)\cdot log(1-p)]$
where $y$ stands for the label and $p$ for probability.
Matthews-Correlation-Coefficient:
One metric is often not enough to fully comprehend the performance of a model and if it is even suited for binary classification problems. Matthews-Correlation-Coefficient (MCC) is becoming more popular in recent years, although it was introduced in 1975 (Matthews, 1975) – it is defined as:
(5) $MCC=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)\cdot(TP+FN)\cdot(TN+FP)\cdot(TN+FN)}}$
However, it is rarely used in web privacy research compared to other fields like web security research. MCC ranges from -1 to +1 and can be understood as a correlation coefficient between the true and predicted label. One particular advantage is the inclusion of values from the confusion matrix, i.e. the MCC will only achieve a high value if both class labels were predicted with high accuracy – which in turn presents the MCC as a meaningful metric for imbalanced datasets (Boughorbel et al., 2017). This was further substantiated by Chicco et al. (Chicco and Jurman, 2020) in 2020, and the authors outlined how accuracy and F1-Score were not reliable measures for imbalanced datasets.
ROC-AUC:
Our next metric is the Receiver Operating Characteristics (ROC) curve, complemented by the Area Under the curve (AUC) value. The former describes a classifier’s performance at various thresholds by plotting the TPR on the ordinate and FPR on the abscissa. The latter metric expresses how well the classifier is able to separate between both class labels.
AUPRC:
The last metric in our study is the Area Under the Precision-Recall Curve (AUPRC). This metric is particularly relevant in the context of a supervised binary classification problem, especially when dealing with imbalanced datasets. The Precision-Recall curve is a plot that illustrates the trade-off between the precision (positive predictive value) and the recall (true positive rate) for a classifier at various thresholds. Precision is defined as the ratio of true positives to the sum of true and false positives, while recall is the ratio of true positives to the sum of true positives and false negatives. The AUPRC is an aggregated measure of this curve, representing the classifier’s ability to correctly identify positive instances across different thresholds. In imbalanced settings, where one class is significantly underrepresented, the AUPRC is a more informative and reliable metric than ROC-AUC, as it focuses on the performance of the classifier on the minority class, which is often of greater interest (Saito and Rehmsmeier, 2015).
6. Experiments
Model Accuracy Log-Loss ROC-AUC AUPRC BACC F1-Score Precision Recall MCC
LR 0.892 0.252 0.86 0.902 0.86 0.812 0.847 0.78 0.738
[0.891;0.894] [0.25;0.255] [0.858;0.862] [0.9;0.904] [0.858;0.862] [0.81;0.815] [0.843;0.85] [0.777;0.784] [0.735;0.742]
GNB 0.853 2.47 0.832 0.792 0.832 0.759 0.74 0.779 0.654
[0.851;0.854] [2.436;2.514] [0.829;0.833] [0.788;0.797] [0.829;0.833] [0.756;0.762] [0.737;0.745] [0.776;0.782] [0.65;0.657]
BNB 0.851 0.617 0.833 0.849 0.833 0.759 0.732 0.789 0.652
[0.849;0.852] [0.608;0.627] [0.83;0.835] [0.847;0.852] [0.83;0.835] [0.757;0.763] [0.729;0.736] [0.784;0.793] [0.648;0.657]
DT 0.956 0.546 0.947 0.943 0.947 0.926 0.928 0.925 0.895
[0.955;0.957] [0.531;0.566] [0.946;0.948] [0.94;0.945] [0.946;0.948] [0.925;0.928] [0.926;0.93] [0.922;0.927] [0.893;0.897]
RF 0.96 0.161 0.951 0.978 0.951 0.932 0.935 0.929 0.903
[0.959;0.96] [0.155;0.168] [0.949;0.952] [0.977;0.979] [0.949;0.952] [0.93;0.934] [0.933;0.937] [0.926;0.931] [0.901;0.906]
ET 0.964 0.166 0.953 0.98 0.953 0.938 0.952 0.925 0.913
[0.963;0.965] [0.159;0.174] [0.952;0.954] [0.978;0.98] [0.952;0.954] [0.937;0.94] [0.95;0.954] [0.923;0.927] [0.911;0.915]
AdaBoost 0.888 0.574 0.856 0.894 0.856 0.806 0.838 0.776 0.729
[0.887;0.89] [0.574;0.575] [0.854;0.858] [0.892;0.897] [0.854;0.858] [0.803;0.809] [0.834;0.841] [0.772;0.78] [0.725;0.733]
GBM 0.895 0.244 0.867 0.916 0.867 0.82 0.845 0.795 0.747
[0.894;0.897] [0.242;0.246] [0.865;0.869] [0.914;0.918] [0.865;0.869] [0.817;0.822] [0.842;0.848] [0.792;0.799] [0.743;0.75]
LGBM 0.931 0.174 0.909 0.957 0.909 0.88 0.908 0.854 0.833
[0.93;0.932] [0.172;0.175] [0.907;0.91] [0.956;0.958] [0.907;0.91] [0.878;0.883] [0.906;0.911] [0.851;0.857] [0.83;0.836]
HistGB 0.929 0.175 0.907 0.956 0.907 0.877 0.902 0.853 0.827
[0.927;0.93] [0.173;0.176] [0.905;0.909] [0.955;0.957] [0.905;0.909] [0.875;0.879] [0.9;0.905] [0.85;0.856] [0.825;0.831]
XGBoost 0.928 0.172 0.912 0.956 0.912 0.878 0.885 0.872 0.827
[0.927;0.93] [0.17;0.174] [0.91;0.914] [0.955;0.957] [0.91;0.914] [0.877;0.881] [0.882;0.889] [0.869;0.875] [0.825;0.831]
XGBoost 0.883 - - - - 0.867 0.867 0.867 -
RF 0.881 - - - - 0.866 0.862 0.871 -
GBM 0.858 - - - - 0.844 0.836 0.856 -
SVC 0.848 - - - - 0.828 0.819 0.852 -
AdaBoost 0.840 - - - - 0.824 0.816 0.836 -
DT 0.834 - - - - 0.814 0.810 0.819 -
KNN 0.833 - - - - 0.814 0.809 0.820 -
LR 0.809 - - - - 0.796 0.789 0.822 -
Table 4. Performance comparison of the $\text{Chrome}_{\textbf{22}}$-trained classifiers. Best results are bold, second best are underlined, and baseline models are in italic. The performance results of the t.ex-Graph classifier are reported in the bottom half (Raschke et al., 2023).
Figure 5. The radar charts illustrate that while all models perform well on the training set, the XGBoost model exhibits more consistent performance across the various test sets, suggesting better generalizability. The other models show reduced performance on test sets, indicating potential overfitting to the training data.
In this section, we evaluate the effectiveness of our proposed methodology to detect trackers. We start by presenting our classifier results based on $\text{Chrome}_{22}$ data and follow-up by testing them with data from $\text{Firefox}_{22}$ and $\text{Brave}_{22}$. Furthermore, we assess the longitudinal performance with the $\text{Chrome}_{23}$ dataset that was collected almost a year later.
We test our approach on an Apple computer, namely MacBook Pro (14-inch, 2021) with a 10-core Central Processing Unit (CPU), a 16-core Graphical Processing Unit (GPU), 1 Terabyte (TB) of SSD storage, and 32 Gigabyte (GB) of Random Access Memory (RAM).
6.1. Intra-browser Performance
Table 4 shows the performance of our $\text{Chrome}_{\textbf{22}}$-trained classifiers and the baselines. In terms of Accuracy, Ensemble Trees (ET) lead with a score of 0.964, closely followed by Random Forest (RF) and Decision Trees (DT) with scores of 0.960 and 0.956, respectively.
The Log-Loss metric, which penalizes classifiers for incorrect predictions, shows that ET, RF, and DT also perform well, with ET having the lowest log loss, indicating a high level of confidence in its predictions. On the other hand, Gaussian Naive Bayes (GNB) exhibits a significantly higher log loss, which suggests that it is less confident in its predictions compared to the tree-based methods. The ROC-AUC metric further strengthens the position of ET, RF, and DT, as they all score above 0.94, indicating that they are good at differentiating between the positive and negative classes. The high MCC scores corroborate the findings from the other metrics, placing these three classifiers as the most robust performers across the board. In contrast, baseline models like LR are comparatively lower performance across these metrics.
The gradient boosting models, particularly XGBoost, have fairly high scores as reflected in an AUC-ROC score of 0.915 and an F1-Score of 0.875. They are able to handle the class imbalance due to the fairly high MCC and AUPRC scores. However, their precision and recall scores indicate room for improvement in identifying instances of the minority class.
Comparing the best performing classifiers ET and RF to the baseline models, shows that they perform significantly higher, except for the DT and its Log-Loss score. The t.ex-Graph classifiers display lower performance across all metrics.
6.2. Cross-browser and Longitudinal Performance
Figure 5 presents the cross-browser performance and longitudinal performance of our $\text{Chrome}_{\textbf{22}}$-trained classifiers. We focus on three selected classifiers, one baseline model, one ensemble tree-based model, and one ensemble tree-based model with gradient boosting.
For Logistic Regression and Random Forest, $\text{Chrome}_{\textbf{22}}$ shows high performance across most metrics, which is expected as the models would be most adapted to the data they were trained on. The performance on $\text{Chrome}_{\textbf{23}}$, $\text{Firefox}_{\textbf{22}}$, and $\text{Brave}_{\textbf{22}}$ varies, with a general trend of decreasing performance, indicating potential overfitting to the training data or a lack of generalizability to other datasets.
The XGBoost model shows a similar trend, with high performance on the training set and slightly varied results on the test sets. However, the variations in the XGBoost model’s performance are less pronounced than in the Logistic Regression and Random Forest models, suggesting that XGBoost may be more robust or generalizable across different browser datasets.
Overall, the figure indicates that while all models perform well on the dataset they were trained on, their ability to generalize to other data is varied. XGBoost appears to have the most consistent performance across different browser datasets, indicating a potential advantage in terms of generalization. The figure underscores the importance of cross-browser testing to ensure that models are not only tuned to a specific dataset but also maintain their performance across diverse scenarios.
The Brave results show a high accuracy metric for the overall performance, but the F1-Score and MCC are low and represent a more realistic performance of the classifiers, which is expected due to the more considerable differences between the Chrome and Brave data.
The longitudinal performance analysis, as depicted in Figure 5, indicates a slight decline in the efficacy of the classification models when transitioning from the $\text{Chrome}_{\textbf{22}}$ to the $\text{Chrome}_{\textbf{23}}$ dataset. This is evidenced by the contraction of the metric lines from $\text{Chrome}_{\textbf{22}}$ to $\text{Chrome}_{\textbf{23}}$ across the three models, suggesting a decrease in metrics such as AUC-ROC, Accuracy, and others. Notably, the XGBoost model demonstrates the smallest degree of performance reduction, underscoring its robustness and adaptability to changes over time. In contrast, both the Logistic Regression and Random Forest models exhibit a more pronounced drop, highlighting potential challenges in model generalization across browser version updates. This suggests that while all models are affected by evolving data characteristics, ensemble methods, particularly gradient boosting as implemented in XGBoost, may offer a more stable predictive performance in longitudinal settings. Applying isotonic calibration had only minor improvements compared to the uncalibrated models as shown in Figure 6.
7. Discussion
Tracking affects users across all three browsers. However, this varies to a certain degree, e.g. Brave had the least amount of tracking-related HTTP responses compared to Chrome and Firefox.
Nevertheless, an imbalance can be observed in all data sets, which might be a problem for specific ML models. Even though many different methods were developed to balance data sets, due to the moderate imbalance in Chrome and Firefox and the good performance, a resampling was not conducted to preserve the observed distribution. Still, the large number of observations provides sufficient HTTP response messages for each class label. Furthermore, several evaluation metrics were calculated and provided that are aimed at imbalanced data sets, e.g. BACC.
The classifiers performed similarly in Chrome and Firefox but were poor with Brave data. First and foremost, the application of binary-encoded features resulted in surprisingly good results, i.e. the mere presence of a header value is enough and extensive feature engineering is not required from a pure tracking detection standpoint. On the other hand, explaining the models is limited, except for model parameters, training time, or the calculated metrics. In addition, this method simplifies the models to a high degree, although HTTP header values can reveal more information about trackers and their differences from non-trackers. Nevertheless, the binary approach helped preserve the core information, namely whether a header is related to tracking.
The NB classifier did not perform as well as others, but that might be due to its limited applicability to the data sets. As mentioned above, NB is based on Bayes Theorem and assumes that features within a class are not correlated to each other. It can be assumed that a set of headers and thus features are correlated to a certain degree. These were not removed which is a violation of Bayes Theorem in a statistical sense could be argued, and thus, its applicability. On the contrary, a paper by Berrar (Berrar, 2018) argues that NB can perform well, even under these circumstances. Other implementations, e.g., the Categorical Naïve Bayes might have been a good fit for the training sets with categorical features (nb, 2023).
Filter lists are reliable and effective when it comes to blocking trackers (Mazel et al., 2019), and are one of the oldest methods, e.g. EasyList was created in 2005 (eas, 2023a). They present a popular ground truth to label data sets for training ML models. However, they have to be constantly updated, which commonly happens manually. A longitudinal study of nine years by Alrizah et al. (Alrizah et al., 2019) showed that human errors, e.g. acting in a timely manner or incorrect reports of possible new trackers and ads, are a root cause for lower performance. Chen et al. (Chen et al., 2021) highlighted evasion methods to circumvent blockers and filter lists: (i) changing URLs, (ii) embedding code directly into the website with the script tag, and (iii) combining functional and tracking code in one file. In summary, using filter lists for our ground truth is not without problems and directly affect the performance of our classifiers.
7.1. Limitations
Firstly, the analysis of response headers to fill the identified research gap has its strengths as well as its weaknesses. Headers represent only a part of an HTTP message, and research has shown that HTTP bodies are also relevant when it comes to understanding and identifying trackers (Li et al., 2015; Guarino. et al., 2020). From a broader point of view, only one part of application layer traffic is analyzed, disregarding other parts of web traffic, e.g. transport layer data.
Secondly, the data set could have been processed differently across the pipeline. As above-mentioned, the data sets are imbalanced, with a 72:28 non-tracker-to-tracker ratio for Chrome and Firefox data. It is not a severe form of imbalance, but it might be problematic for some classification algorithms. They could be biased against the majority class, and it might be possible to consistently predict the non-tracker label and still observe good metrics. Several resampling techniques exist, simple methods, e.g. over- and undersampling (Hernandez et al., 2013), or more complex methods, e.g. SMOTE (Chawla et al., 2002), or one of its variations (Han et al., 2005; Mukherjee and Khushi, 2021). Additional measures include using ML models, which handle imbalanced data sets well (Juez-Gil et al., 2021), or anomaly detection (Park et al., 2018). In summary, this aspect of the data pipeline has not been implemented with any of the aforementioned methods, and it might be interesting to explore how well these techniques may work for the present concept.
Thirdly, the labeling of the initial data sets, i.e. the quality of the ground truth. T.EX allows the implementation of different filter lists and offers three to date, namely EasyList, EasyPrivacy, and Disconnect.me. All base their functionality on extensive filter lists, i.e. rules and RegEx are used to block tracking-related HTTP requests. Updates are carried out regularly every four days for the first two (eas, 2023a, b), and on an irregular schedule for the latter (dis, 2023). Research has shown that these filter lists are not without errors regarding their rules (Snyder et al., 2020).
8. Conclusion
In this work, we trained Chrome-based classifiers on a subset of binary-encoded HTTP/S response headers to evaluate their effectiveness and usefulness in web tracker detection while retaining the natural imbalance of trackers. In addition, we measured the cross-browser effectiveness, showing that multiple Chrome-based classifiers achieve high accuracies for Firefox as well. However, large performance losses were observed for Brave. The longitudinal performance of the selected ML models deteriorated only by small margins, which highlights their potential due to the changing capabilities of trackers (Fouad et al., 2018; Bujlow et al., 2017; Englehardt and Narayanan, 2016). Response headers can be used to differentiate well between trackers and non-trackers and they provide additional information to the nature of tracking-related entities.
For our future research we plan to further investigate response headers, focusing on the values for feature engineering, or the use of Deep Learning models. Lastly, we will approach the deployment aspect two-fold, one from a user perspective in form of browser extension to test the models in a real-life scenario, and secondly as a fully automated system in the cloud that integrates a crawler to periodically update the models.

### v2

@@
 The World Wide Web’s connectivity is greatly attributed to the HTTP protocol, with HTTP messages offering informative header fields that appeal to disciplines like web security and privacy, especially concerning web tracking.
-Despite existing research employing HTTP/S request messages to identify web trackers, HTTP/S response headers are often overlooked.
-This study endeavors to design effective machine learning classifiers for web tracker detection using HTTP/S response headers.
+Despite existing research employing HTTP request messages to identify web trackers, HTTP response headers are often overlooked.
+This study endeavors to design effective machine learning classifiers for web tracker detection using binarized HTTP response headers.
 Data from the Chrome, Firefox, and Brave browsers, obtained through the traffic monitoring browser extension T.EX, serves as our dataset.
-Eleven supervised models were trained on Chrome data and tested across all browsers.
+Ten supervised models were trained on Chrome data and tested across all browsers, including a Chrome dataset from a year later.
 The results demonstrated high accuracy, F1-score, precision, recall, and minimal log-loss error for Chrome and Firefox, but subpar performance on Brave, potentially due to its distinct data distribution and feature set.
-The research suggests that these classifiers are viable for detecting web trackers in Chrome and Firefox.
+The research suggests that these classifiers are viable for web tracker detection.
 However, real-world application testing remains pending, and the distinction between tracker types and broader label sources could be explored in future studies.
-Web Privacy, Web Tracking, Web Privacy Measurement, HTTP/S Response Header, Machine Learning
+Web Privacy, Web Tracking, Web Privacy Measurement, HTTP Response Header, Machine Learning
 journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
@@
 The rapid evolution of the Web has cemented its role as an integral component of modern life.
-Through desktop and mobile browsers, users access a vast array of information and services.
+Users access a vast array of information and services through desktop and mobile browsers.
 However, every interaction, from content consumption to mere site visitation, reveals information about the habits and preferences of users, which in turn has implications for personal privacy and online security.
 These privacy concerns are highlighted by the ubiquity of behavioural targeting.
-By analyzing a user’s browsing history, a detailed profile can be constructed, facilitating the delivery of tailored advertisements and content.
+A detailed behavioural profile can be constructed by analyzing a user’s browsing history, facilitating the delivery of tailored advertisements and content.
 This profiling is made possible by web tracking that employs sophisticated techniques to track users across the Web.
 Given the increasing concerns about such practices (Melicher et al., 2016; Chanchary and Chiasson, 2015), there is a growing emphasis on research in web privacy.
-An established countermeasure against unsolicited HTTP communications is the use of filter lists (Traverso et al., 2017), such as EasyList (eas, 2023a) and EasyPrivacy (eas, 2023b).
+An established countermeasure against unsolicited HTTP communications is the use of filter lists (Traverso et al., 2017) such as EasyList  [^1] https://easylist.to/easylist/easylist.txt (EL) and EasyPrivacy  [^2] https://easylist.to/easylist/easyprivacy.txt (EP).
 However, their dependency on manual curation makes them vulnerable to oversights and rapid obsolescence (Hashmi et al., 2019; Alrizah et al., 2019; Fouad et al., 2018).
 Moreover, some trackers employ strategies like domain masking to evade these lists (Chen et al., 2021).
-In an attempt to mitigate these shortcomings, recent efforts have explored the potential of Machine Learning (ML) in this domain.
-Various classifiers, some achieving accuracy metrics in the 80th percentile (Cozza et al., 2020; Iqbal et al., 2021), have been developed using diverse data sources such as HTTP headers and JavaScript calls.
-Yet, a majority of these focused primarily on HTTP requests, potentially neglecting other informative data streams.
-Additionally, the predominant reliance on Firefox for data collection may limit the comprehensiveness of these studies.
-Notably, there seems to be a gap in literature regarding the use of HTTP response headers as the sole data source for tracker classification across diverse browsers, despite studies in the related field of online security highlighting the efficacy of this approach (McGahagan et al., 2019).
-This study intends to bridge this gap.
-We strive to develop classifiers adept at web tracker detection, while confronting challenges such as data sparsity, high dimensionality, and class imbalance.
-Our key contributions are:
+In an attempt to mitigate these shortcomings, efforts of past decades have explored the potential of Machine Learning (ML) in this domain.
+Various classifiers, some achieving accuracy metrics in the 80th percentile (Cozza et al., 2020; Iqbal et al., 2021), have been developed using diverse data sources such as HTTP headers and JavaScript API calls.
+Meanwhile, the potential of information contained in HTTP responses has been largely overlooked, except for HTTP cookies.
+The value of this information might be doubted because it becomes available only after web tracking has occurred.
+However, this limitation primarily affects scenarios where the model’s predictions are used to block communications with web trackers in real-time.
+Alternative approaches are feasible: (i) dynamically generating filter lists based on the model’s predictions, or (ii) using it as a supplementary tool for classifiers that depend solely on pre-HTTP request information.
+Most previously proposed classifiers have not addressed the issue of deployability, suggesting that the field is still in an exploratory phase.
+Consequently, it is valuable to examine the suitability of HTTP response information for web tracker classification.
+Unlike HTTP request information, which is controlled by the client and varies between users, HTTP response information is configured by the server, i.e., the potential web tracker.
+Notably, a study in the related field of web security underscores the effectiveness of this approach (McGahagan et al., 2019).
+Our key contributions are as follows:
   1. (1)
-    We conduct a rigorous study of HTTP response headers across multiple browsers, focusing on their potential for web tracker detection within the Tranco top 10k websites.
-Furthermore, we evaluate dataset similarities to discern differences in HTTP response headers across the assessed browsers.
+    We conduct a study of HTTP response headers across multiple browsers, focusing on their potential for web tracker detection within the Tranco global top 10K websites.
+Furthermore, we identify challenges with this type of data for developing ML classifiers.
   2. (2)
-    We present a semi-automated ML pipeline which facilitates the training of eleven classifiers tailed for web tracker detection.
-Our findings indicate that selected tree and gradient boosting models, when applied to a subset of binary-encoded HTTP/S headers, are capable of achieving ROC-AUC, AUPCR, and F1 scores exceeding $0.9$.
+    We present a semi-automated ML pipeline which facilitates the training of ML models for the binary classification of web trackers based on HTTP response or request headers.
   3. (3)
-    We conduct a comprehensive and multi-faceted evaluation of these classifiers, which includes longitudinal studies to gauge potential concept-drift and evaluates the cross-browser performance of classifiers trained solely on traffic data collected within the Chrome browser.
-Notably, our evaluation incorporates thirteen metrics, offering a holistic view of the binary classification challenge.
+    We train and evaluate ten supervised classifiers.
+Our findings indicate that when applied to a subset of binarized HTTP headers, selected tree and gradient boosting models can detect web trackers.
+Our best classifier achieves ROC-AUC, AUPRC, and an F1-score exceeding $0.93$.
   4. (4)
-    We assess calibration effects on final classifier performance — a consideration seldom addressed in prior research.
-This allows us to discern if models trained solely on data from one browser can be effectively deployed on data from alternate browsers without the necessity for retraining.
-The remainder of this paper is structured as follows.
-Section 2 reviews past research in the field of web tracker detection.
+    We conduct a comprehensive and multi-faceted evaluation of these classifiers, which includes longitudinal studies to gauge potential concept drift and evaluates the cross-browser performance of classifiers trained solely on traffic data collected within the Chrome browser.
+We show that Chrome data captures a majority of the structural header characteristics of trackers.
+Our evaluation incorporates 13 metrics, offering a holistic view of the binary classification challenge.
+In addition, we use our pipeline to train classifiers using only HTTP request headers to highlight the comparative value of HTTP response headers for web tracker detection.
+The remainder of this paper is structured as follows: Section 2 reviews past research in the field of web tracker detection.
 Section 3 introduced definitions, our research questions, and the analyzed datasets.
-Section 4 explores HTTP/S response headers from Chrome, Firefox, and Brave data sets, providing general findings and identifies potential challenges for the ML pipeline.
+Section 4 explores HTTP response headers from Chrome, Firefox, and Brave datasets, providing general findings and identifying potential challenges for the ML pipeline.
 Section 5 describes the design and implementation of the ML pipeline, including selected models, metrics, and baselines.
 Section 6 presents the evaluation of the trained classifiers.
-The penultimate Section 7 discusses our results and limitations.
-Section 8 concludes the paper.
+The penultimate Section 7 discusses our results and limitations, and Section 8 concludes the paper.
 2. Related Work
@@
 2.1. Classification Based on HTTP Information
-Approaches to the detection web trackers with the support of machine learning date back to the early 2000s.
+Approaches to the detection of web trackers with the support of machine learning date back to the early 2000s.
 In 2010, Yamada et al.
-(Yamada et al., 2010) were the first to use temporal link analysis to build a graph based on hosts related to HTTP requests from the traffic generated by corporate network.
-Their classifier considers temporal information such as the visiting time of a specific website or host.
+(Yamada et al., 2010) were the first to use temporal link analysis to build a graph based on hosts related to HTTP requests from the traffic generated by corporate networks.
+Their classifier considers temporal information, such as the visit duration at a specific website or host.
 Yamada et al.
@@
 (Bau et al., 2013) described and discussed the usefulness of machine learning as an effective counter to web tracking.
-They outlined the problems with manually curated filter lists - problems that still hold true today, such as the reliance on manual curation and the inflexibility of static lists.
+They outlined problems with manually curated filter lists - problems that still hold true today, such as the reliance on manual curation and the inflexibility of static lists.
 Beyond building an Elastic Net model focused on traces of content extraction in the Document Object Model and HTTP headers to show the advantages of automated approaches, Bau et al.
@@
 In 2015, Gugelmann et al.
-(Gugelmann et al., 2015) proposed an alternative approach by deriving several statistical features from HTTP traffic, e.g.
-up- and download volume.
+(Gugelmann et al., 2015) proposed an alternative approach by deriving several statistical features from HTTP traffic, e.g., up- and download volume.
 In contrast to prior studies, they thoroughly described their model-building process and emphasized their feature engineering.
@@
 (Li et al., 2015) proposed TrackAdvisor, a system to classify HTTP requests based on the collective statistics of cookies contained therein.
-Using FourthParty, they captured 563,031 requests and identified 22,270 cookies that were sent to third parties.
-Rather than relying solely on filter lists, they used a multi-step process with an emphasis on manual inspection to label the captured data.
-Notably, the training and test sets created from the captured data comprised only 500 requests each and had a balanced class distribution, which contrasts with the distribution found in the original data.
+Using FourthParty, they captured 563,031 requests and identified 22,270 cookies sent to third parties.
+Rather than relying solely on filter lists, they used a multi-step process emphasising manual inspection to label the captured data.
+Notably, the training and test sets created from the captured data comprised just 500 requests each and had a balanced class distribution, which differs from the distribution found in the original data.
 Nevertheless, Li et al.
-used these sets to train a SVM that ultimately achieved a precision of 99.4% and a recall of 100%.
+used these sets to train an SVM that ultimately achieved a precision of 99.4% and a recall of 100%.
 Like Gugelmann et al., they point out the usefulness of the HTTP Referrer header for third-party tracker detection.
 In 2016, Dudykevych and Nechypor (Dudykevych and Nechypor, 2016) also used cookies as the basis for their classifiers.
-In contrast to Li et al., however, they developed their own crawler, had a larger data set of 550 000 cookies, and developed a different feature set.
-The latter is only described briefly, but information from the cookies themselves and the HTTP requests were used as features.
-It can be assumed that the features are numerical and categorical, as the authors used Principal Component Analysis and correlation, which require numerical data, as well as binary encoding of categorical features.
+In contrast to Li et al., however, they developed their own crawler, used a larger dataset of 550,000 cookies, and developed a different feature set.
+The latter is only briefly described, but information from the cookies and HTTP requests were used as features.
+It can be assumed that the features are numerical and categorical, as the authors used Principal Component Analysis and correlation, which require numerical data, as well as binarization of categorical features.
 The resulting classification accuracies ranged from 82.01% for Naïve Bayes to 95.43% for the regularized logistic regression method.
-The analysis of HTTP traffic for and with machine learning models is also an important topic in web security research, albeit with a different target (i.e.
-label) than web privacy research.
+The analysis of HTTP traffic for and with machine learning models is also an important topic in web security research, albeit with a different target (i.e., label) than web privacy research.
 In a 2019 study, McGahagan et al.
 (McGahagan et al., 2019) analyzed response headers from 45 874 websites and extracted 672 features, resulting in a final feature vector with 22 properties.
-In addition, they tested whether over- and undersampling had any effects on the model’s performance.
-The coverage of metrics, models, different training and tuning combinations, and extensive feature engineering documentation can be considered standard among machine-learning practitioners but is somewhat rare among prior related research.
-An important contribution of this paper is the in-depth presentation of the usefulness of HTTP response headers for automated ML models.
+In addition, they tested whether over- and undersampling affected the model’s performance.
+Their extensive coverage of metrics, models, different training and tuning combinations, and feature engineering documentation can be considered standard among ML practitioners but is somewhat rare among related research in the field of web tracking detection.
+Crucially, an important contribution of this paper is the in-depth presentation of the usefulness of HTTP response headers for ML models to detect malicious entities.
+We apply a similar approach but to web privacy research.
 Unlike most papers discussed in this section, which have used data from the Firefox browser to test the performance of supervised learning models, Din et al.
@@
 (Cozza et al., 2020) developed a system that not only classifies HTTP requests but also automatically updates the filter list.
-Although this study is not the first to detect previously undocumented trackers using machine learning techniques, the deployment of a tracker-detecting browser extension that combines a filter list with a machine learning models and updates its ground truth marks an important step forward in tracker detection technology.
+Although this study is not the first to detect previously undocumented trackers using machine learning techniques, the deployment of a tracker-detecting browser extension that combines a filter list with a machine learning model and updates its ground truth marks an important step forward in tracker detection technology.
 Guarino et al.
@@
 and Li et al.
-by training multiple model based on combinations of features used by these studies, including different HTTP headers and the ratio of upstream and downstream data transmissions.
+by training multiple models based on combinations of features used by these studies, including different HTTP headers and the ratio of upstream and downstream data transmissions.
 Their top performer is a Random Forest model with an average accuracy of 92%.
@@
 (Iqbal et al., 2022) conducted an extensive study that uses entire request chains rather than individual requests as the basis for classifying web tracking.
-Notably, they were among the first to consider responses as well as requests by developing three distinct feature sets: the first is based on the entire request chain, the second relies on information pertaining to individual requests and the third considers response headers and metadata.
+Notably, they were among the first to consider responses alongside requests by developing three distinct feature sets: the first is based on the entire request chain, the second relies on information pertaining to individual requests, and the third considers response headers and metadata.
 By training a Random Forest model on five distinct datasets with different configurations from separate crawls, Iqbal et al.
 achieved accuracies above 98%.
-However, if the model was trained on only one crawl and tested on the others, the accuracy dropped by up to 14%, highlighting the difficulties of applying models to datasets with varying configurations, e.g.
-data collected across different browsers.
+However, if the model was trained on only one crawl and tested on the others, the accuracy dropped by up to 14%, highlighting the difficulties of applying models to datasets with varying configurations, e.g., data collected across different browsers.
+Whereas Iqbal et al.
+selected features based on intuition and domain knowledge, our approach selects features from a data-driven standpoint and considers all response headers for feature extraction.
 2.2. Other Detection Approaches
-Beyond the HTTP-focused approaches outlined above, a variety of classification methods utilizing different data sources for the extraction of distinctive criteria have been proposed and evaluated.
-These include, classifiers focused on JavaScript code and API access events that aim to identify device fingerprinting techniques, pure URL-based classifiers, CNAME cloaking detectors, and graph-based classifiers that leverage the topology of the Web to identify third-party tracking in particular.
-In the following paragraphs, we will briefly highlight notable research on these alternative approaches.
+Beyond the HTTP-focused approaches outlined above, various classification methods utilizing different data sources to extract distinctive criteria have been proposed and evaluated.
+These include classifiers focused on JavaScript code and API access events that aim to identify device fingerprinting techniques, pure URL-based classifiers, CNAME cloaking detectors, and graph-based classifiers that leverage the topology of the Web to identify third-party tracking in particular.
+The following paragraphs briefly highlight notable research on these alternative approaches.
 Focusing on JavaScript API calls, Wu et al.
 (Wu et al., 2016) developed DMTTrackerDetector, a system using decision trees in which observed API calls serve as features.
-The system, comprising a 717-dimensionial feature vector, achieved 97.8% accuracy and addressed obfuscation issues, showing the distinct JavaScript API usage by trackers and non-trackers.
+Their system, comprising a 717-dimensional feature vector, achieved 97.8% accuracy and addressed obfuscation issues, showing the distinct JavaScript API usage by trackers and non-trackers.
 Bahrami et al.
 (Bahrami et al., 2021) refined the scope of their analysis by focusing on device fingerprinting rather than API calls in general.
-Their approach combines graph analysis with supervised and unsupervised learning to detect web trackers in historical JavaScript file data collected between 2010 and 2019 by creating a graph that includes temporal aspects and the output of AST-based keyword extraction.
+Their approach combines graph analysis with supervised and unsupervised learning to detect web trackers in historical JavaScript file data collected between 2010 and 2019 by creating a graph that includes temporal aspects and the output of Abstract Syntax Tree-based keyword extraction.
 Their method, which includes nine graph-related features and graph embedding, as well as a clustering model, yielded accuracies up to 89.13% and identified distinct clusters for fingerprinting and functional APIs, highlighting the viability of such an approach.
@@
 (Castell-Uroz et al., 2021) proposed a tripartite network graph to identify third-party tracking resources and their origin domains within first-party websites.
-By applying this approach to a dataset comprising over 90 thousand domains, they were able to detected new trackers through their hash value popularity and dirt level, finding that higher values correlated with the likelihood of a given resource to be tracking-related.
+By applying this approach to a dataset comprising over 90,000 domains, they detected new trackers through their hash value popularity and dirt level, finding that higher values correlated with the likelihood of a given resource to be tracking-related.
 Conversely, Metwalley et al.
@@
 Their method, applied to 200 Alexa-ranked websites, identified 34 new trackers and successfully supported the analysis of cookie-matching practices, though it had limitations in detecting all known trackers.
-Finally, Yu et al.
-(Yu et al., 2016) proposed a browser extension using k-Anonymity principles to classify third-party trackers based on data sent to users.
-Leveraging data from 200,000 German users of the Cliqz browser, their system outperformed traditional blockers like Disconnect, which had a 38.8% false positive rate.
 2.3. Summary
-Despite over a decade of dedicated research, web tracking is as prevalent as ever, new tracking methods are developed, and the most widespread form of protection, namely blockers and their filter lists, exhibit several key problems.
-Filter lists, which are also used as the ground truth for labeling in many studies related to web privacy, are often maintained by a small group of users, contain irrelevant rules, are rather slow to add new rules, and can simply be circumvented by changing URLs.
-In response, researchers have developed a wealth of automated approaches, utilizing HTTP messages, JavaScript, or the HTML files of websites themselves to classify third parties as trackers.
-A crucial component of this ongoing research effort is the availability and quality of data on which to train and evaluate machine learning approaches.
-Despite this, HTTP response headers are often neglected in this regard.
-There are cases where authors document the collection of HTTP responses (Acar et al., 2013) but not their actual use.
-In other cases, the data is collected, but only used in support of other approaches such as the classification of cookies (Lerner et al., 2016; Englehardt et al., 2015).
-Additionally, the majority of studies are based on datasets collected using just one browser, most commonly Firefox.
-Cassel et al.
-(Cassel et al., 2022) discuss the need for cross-browser examinations to avoid biased datasets.
-This is supported by research that shows that characteristics such as JavaScript API calls differ between browsers (Yang and Yue, 2020).
-In summary, our review of related research has shown that, although the application of machine learning to the field of web tracking identification and classification has been investigated in numerous studies, a research gap still exists with regard to cross-browser performance of classifiers and the usefulness of HTTP response data for classification purposes.
+Research presented in this section explored various approaches to detect web trackers using machine learning, utilizing various available information.
+However, HTTP response information has been mostly neglected, and if considered, only the cookie information or a small subset of header information was taken into account.
+This gap may stem from the intuition that the model’s predictions should be utilized in a real-time application scenario, ignoring use cases where ex-post detection of trackers is sufficient.
+However, prior studies often owe a thorough discussion on the deployability of their approaches.
+Related to this, training and test data have been generated by the same browser, often in the same browsing session.
+The generalization of a model’s predictions is yet to be evaluated.
+Furthermore, the metrics used to assess a classifier’s performance are not homogenized in the related literature, making a direct comparison of the classifiers’ performances unfeasible.
 3. Approach
-In this study, we address the research gap identified in our review of related research by evaluating the (cross-browser) performance of commonly used machine learning classifiers that are trained on HTTP/S response headers.
-In this section, we discuss our underlying research questions and present the approach we used to address them.
+This study addresses the research gap identified in our review of related research by evaluating the (cross-browser) performance of commonly used machine learning classifiers that are trained on HTTP response headers.
+In this section, we present our underlying research questions and the approach we employ to address them.
 3.1. Preliminaries
-A core principle of automated web tracker detection is the definition of what constitutes a tracker.
-The focus hereby lies on the tracking entities and not on the activity of tracking.
-Our definition of what constitutes a tracker or non-tracker is coupled to our ground truth to label each dataset.
-We define them as follows:
+A core principle of automated web tracker detection is delineating what constitutes a tracker.
+The focus of this delineation lies on the tracking entities and not on the tracking activity.
+Our definition of a tracker (T) and non-tracker (NT) is contingent upon the ground truth used to label each dataset:
 […]
@@
 $NT$ =$=\{h\in H\mid\forall r\in R:\text{match}(h,r)=0\}$
-In other words: if the remote hostname (i.e.
-the sender) of a response matches at least one rule in our aggregate list compiled from the EasyList and EasyPrivacy filter lists, we label the response as a tracker.
-For the sake of simplicity, we will use the terms header and response header interchangeably to refer to HTTP/S response headers throughout the remainder of the paper, unless explicitly stated otherwise.
+In other words: if the remote hostname (i.e., the sender) of a response matches at least one rule in our aggregate list compiled from the EL and EP filter lists, we label the response as a tracker.
+For the sake of simplicity, we use the terms header and response header interchangeably to refer to HTTP response headers throughout the remainder of the paper unless explicitly stated otherwise.
 3.2. Research Questions
-The main goal of this study is to analyze whether response headers are useful for automated web tracker detection.
-Based on the ML approaches of prior research and the observed areas to further extend our empirical understanding and detection of web trackers, we formulate the following research questions that guide us throughout this study:
-  1. (1)
-    What are the differences and similarities between web trackers and non-trackers with regard to their response headers?
-  2. (2)
-    Can trained classifiers detect web trackers with only response header features while achieving an F1-Score higher than 0.9 in an imbalanced setting?
-  3. (3)
-    How large is a classifier’s performance degradation when it is trained on one browser’s HTTP traffic data and then applied to other browsers?
-To answer these research questions, we develop a semi-automated ML pipeline that transforms HTTP traffic datasets collected with T.EX (Raschke et al., 2019) and uses it to train several supervised binary classifiers.
-We use reliable data that was compiled by prior research to facilitate future comparisons between new classifiers.
-To test the robustness and reliability of our classifiers, we train them on HTTP data from Chrome and apply them to different datasets with a variety of configurations, including other browsers (latitudinal comparison) as well as different time frames within the same browser (longitudinal comparison).
-Furthermore, past research observed an uneven distribution of trackers and non-trackers, which is why we train our classifiers with an imbalanced distribution, to assess how effective they perform with the likely true distribution.
+The primary objective of this study is to assess the efficacy of response headers for automated web tracker detection.
+Based on the ML approaches of prior research and the observed areas to further extend our empirical understanding and detection of web trackers, we formulate the following research questions (RQ) that guide us throughout this study:
+  • RQ1: What are the key characteristics of HTTP response header data in the web tracking domain? We want to understand the datasets to develop our pipeline and classifiers accordingly.
+  • RQ2: Can trained classifiers detect web trackers with only response header features in an imbalanced setting while achieving a similar performance to past research?
+We aim to train high-performing classifiers while keeping the likely true distribution.
+We define our lower bound as an F1-Score of 0.9 based on the classification results of past research to provide a meaningful addition to existing detection solutions.
+Trackers and non-trackers are not equally present in the datasets, thus an imbalance is present and we want to assess how well classifiers perform in this setting.
+  • RQ3: How large is a classifier’s performance degradation when it is trained on one browser’s HTTP traffic data and then applied to other browsers?
+Web trackers might exhibit different header characteristics depending on the browser.
+We use data from three browsers to capture a potential variety of web trackers and to assess the stability and generalization stability of our classifiers.
+We address our RQs by developing a semi-automated ML pipeline that transforms HTTP traffic datasets collected with T.EX - The Transparency EXtension (T.EX) (Raschke et al., 2019) and use it to train several supervised binary classifiers.
+Additionally, we use reliable data compiled by prior research to facilitate future comparisons between new classifiers.
+Semi-automated means that the ML pipeline has to be started manually, is based on manual data exploration, and does not employ any AutoML libraries.
+The pipeline executes all steps automatically but cannot react to structurally new datasets without any changes or human intervention.
+Fully automating the pipeline requires the development of a direct interface between T.EX and the pipeline, i.e., automatically running new crawls and importing the new datasets to the pipeline, as well as structural converters for datasets from other sources.
+To test the robustness and reliability of our classifiers, we train them on a Chrome dataset and apply them to other browser datasets (latitudinal comparison) and another Chrome dataset generated one year later (longitudinal comparison).
+Furthermore, as past research observed an uneven distribution of trackers and non-trackers, we train our classifiers with the original imbalanced distribution to assess how effectively our classifiers perform on the true distribution.
 $\text{Chrome}_{22}$ $\text{Firefox}_{22}$ $\text{Brave}_{22}$ $\text{Chrome}_{23}$
 T NT T NT T NT T NT
-#Responses 256 202 602 516 309 391 526 865 2969 561 694 206 747 596 066
+#Responses 256,202 602,516 309,391 526,865 2969 561,694 206,747 596,066
 #Responses in % $\approx 0.3$ $\approx 0.7$ $\approx 0.37$ $\approx 0.63$ $\approx 0.5$ $\approx 99.5$ $\approx 0.26$ $\approx 0.74$
@@
 Unique Headers per response ($Q1,\tilde{\mu},Q3$) 10,13,16 12,15,19 12,15,18 13,16,19 13,18,20 12,15,19 10,13,16 13,15,19
-Table 1. Overview of the collected data sets from (zen, 2023), split by data sets as well as tracker (T) and non-trackers (NT).
-Figure 1.
-This figure illustrates the usage headers across four different browsers.
-The left plot (a) displays the frequency distribution of headers for each browser, showing how often each header is present in the dataset.
-The right plot (b) presents a boxen plot of the same data, which provides a more detailed view of the distribution, revealing variations and trends among the browsers.
+Table 1. Overview of the collected datasets from (Raschke, 2023a), split by datasets as well as tracker (T) and non-trackers (NT).
 3.3. Data Collection
-We used existing data sets, containing labeled HTTP response headers by EasyList and EasyPrivacy.
-Following suggested good practices by (Demir et al., 2022), the crawl should fulfill three citeria: (i) repeated crawls across multiple browsers, namely Chrome, Firefox, and Brave, (ii) the documentation of the geographical region as well as the technical environment, and (iii) web sites from the Tranco ranking as it proved to be less biased than the Alexa ranking (Pochat et al., 2018).
-A previously crawled data set by Raschke using T.EX - The Transparency EXtension, fulfills all three criteria and is described as follows:
+We use existing datasets containing labeled HTTP messages by EL and EP.
+Following best practices suggested by (Demir et al., 2022), the crawls that generate the datasets should fulfill three criteria: (i) repeated crawls across multiple browsers, (ii) documentation of the geographical region as well as the technical environment, and (iii) selection of web sites is based on the Tranco ranking, which was proven to be less biased than the Alexa ranking (Pochat et al., 2018).
+An open-source dataset by Raschke that used T.EX fulfills all three criteria and is described as follows:
 Multiple simultaneous and stateful crawls (6 x Chrome, 6 x Brave, 6 x Firefox) of the Tranco top 10K websites (as of 16th of May 2022) were performed on the 12th of August 2022 with T.EX […].
@@
 Note: 2 of the 6 Firefox instances crashed during the crawl.
-Therefore, only 4 Firefox datasets are available (zen, 2023).
-To assess a potential concept drift and therefore model degradation over time, we crawled one more data set from 2023 for Chrome.
-We followed a similar methodology as described in (zen, 2023), using T.EX with Chrome to explore the top 10K websites from the Tranco ranking as of March 29th, 2023.
-Our reasons to choose only Chrome for a new data collection are as follows: Chrome is the most widely used in the real world with a market share of 66.13% on desktops, whereas Firefox has 7.1% (sta, 2023) and Brave has less than 1% (kin, 2023).
-Based on this distribution, the Chrome data set serves as the foundation for training to reflect and reach the majority, i.e.
-using data that most users encounter themselves.
-In addition, web privacy research often relies on a simulated Firefox browser with Selenium due to the use of OpenWPM, leaving space for other browsers of interest.
-In addition, it helps us to answer the question whether a model based on Chrome data can be applied to Firefox or Brave while achieving high performance metrics?.
+Therefore, only 4 Firefox datasets are available (Raschke, 2023a).
+To assess potential concept drift and resulting model degradation over time, we crawled one more dataset for Chrome in 2023 and made it open-source (Raschke, 2023b).
+We follow a similar methodology as described in (Raschke and Cory, 2022), using T.EX and visiting the landing pages with Chrome to explore the top 10K websites from the Tranco ranking as of March 29th, 2023.
+Our reasons for only generating an additional dataset for Chrome are as follows: firstly, Chrome is the most widely used browser with a market share of 66.13% on desktops, whereas Firefox has 7.1%  [^3] https://gs.statcounter.com/browser-market-share/desktop/worldwide and Brave has less than 1%  [^4] https://kinsta.com/browser-market-share/.
+Based on this distribution, the Chrome dataset serves as the foundation for training to reflect and reach the majority, i.e., using data most users encounter.
+Secondly, web privacy research often relies on a simulated Firefox browser with Selenium due to the use of OpenWPM, leaving space for other browsers of interest.
+Thirdly, it helps us answer whether a model based on Chrome data can be applied to Firefox or Brave while achieving high-performance metrics.
 4. Initial Response Header Analysis
-To get a better understanding of the data underlying our analysis before designing our machine learning pipeline, we conduct an initial analysis of the data, the results of which we present in this section.
-We first explore response headers across all three browsers to find empirical evidence of their potential relation to trackers.
-Furthermore, we identify the characteristics of these datasets to apply the appropriate methods in our ML pipeline.
-We use these insights to develop our classifiers to evaluate the effectiveness of response headers for web tracker detection.
+To address RQ1 and gain a better understanding of the data underlying our analysis before designing our ML pipeline, we conduct an initial analysis of the data, the results of which we present in this section.
+Our exploration is guided by intuition and domain knowledge.
+We first explore headers across all three browsers to find empirical evidence of their potential relation to trackers.
+The goal is to identify the key characteristics of the datasets and not to present a complete analysis.
+We then use these insights to develop our classifiers and evaluate the effectiveness of response headers for web tracker detection.
 4.1. Descriptive Analysis of Headers
-A descriptive analysis is useful to gain an understanding of the data structure and the distribution of classes and features within the dataset.
+First, we conduct a descriptive analysis to understand the data structure and distribution of class and feature features within the datasets.
 Table 1 summarizes descriptive metrics for each dataset.
-We start with the datasets $\text{Chrome}_{22}$, $\text{Chrome}_{23}$, and $\text{Firefox}_{22}$.
-All three datasets show similar results for the distribution between trackers and non-trackers, the number of unique headers in total and per response.
-For example, we can observe a small quantitative change in the number of unique headers between $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$.
-$\text{Chrome}_{22}$ lists a total of 7577 unique headers, while $\text{Chrome}_{23}$ has a slightly lower total of 7398 for non-trackers.
-When comparing trackers, $\text{Chrome}_{22}$ had 2213 unique headers, which is nearly the same as $\text{Chrome}_{23}$’s count of 2164.
+We begin with the datasets $\text{Chrome}_{22}$, $\text{Chrome}_{23}$, and $\text{Firefox}_{22}$.
+All three datasets show similar results for the distribution between trackers and non-trackers, as well as the number of unique headers in total and per response.
+Let $B$ represent the set of browsers, where $B=\{\text{Chrome}_{22},\allowbreak\text{Chrome}_{23},\allowbreak\text{Firefox}_{22},\text{Brave}_{22}\}$ and let $L$ represent the set of labels, where $L=\{T,NT\}$.
+For each browser $b\in B$ and each label $l\in L$, we define $HN_{b,l}$ as the set of all distinct header names across all HTTP response messages from a browser $b$ and categorized under label $l$.
+The number of unique headers for each combination of $b$ and $l$ is defined as the cardinality of $HN_{b,l}$:
+$|HN_{b,l}|=|\{hn\mid hn\in HN_{b,l}\}|$
+We can observe a small quantitative change in the number of unique headers between $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$.
 On average, non-trackers have more unique headers as well as a higher median $\tilde{\mu}$ and interquartile range than trackers.
-The $\text{Brave}_{22}$ browser with its standard configuration is the only one that has a low number of trackers and unique headers in tracking-related response messages.
-Interestingly, the $\tilde{\mu}$ equals 18 for trackers and 15 for non-trackers, which represents the only instance where trackers had a higher number.
-Although the dataset has almost only non-trackers, the number of unique headers is similar to the other datasets.
-Figure 1 shows a frequency plot, which is a distribution graph that shows how often each header appears in the dataset.
-It uses a logarithmic scale to manage the wide range of frequencies, which is a common approach when data spans several orders of magnitude.
-This plot typically demonstrates a power-law distribution, characterized by a steep drop-off where a small number of headers are occurring frequently, while the majority are rare.
-This type of distribution is common in HTTP data, where few headers dominate, and many others have a long tail of low frequency.
+As Brave is a privacy-oriented browser that performs extensive blocking of web tracking by default, $\text{Brave}_{22}$ exhibits a far lower number of trackers and unique headers in tracking-related response messages than the other datasets.
+Interestingly, the $\tilde{\mu}$ equals 18 for trackers and 15 for non-trackers, representing the only instance where trackers had a higher number.
+Although the dataset comprises almost exclusively non-trackers, the number of unique headers is similar to the other datasets.
+Figure 1. Frequency distribution of headers for each browser, indicating the frequency with which each header is present in the dataset.
+Figure 1 shows a frequency plot to outline how often each header appears in the dataset.
+The plot demonstrates a power-law distribution, where few headers occur frequently, and many others form a long tail of low frequency.
 This sparseness indicates that while there are many different headers, only a few are consistently used across different websites and by different browsers.
-The boxen plot graph of Figure 1 offers a visual representation of the distribution of header frequencies across different browsers and versions.
-The spread and median values suggest that the distribution of headers is not uniform across browsers, with some browsers having a higher tendency to use certain headers than others.
-Furthermore, if the median line in the boxen plot is closer to the bottom of the box, it suggests a right-skewed distribution, aligning with what the frequency plot shows — a few headers with high frequency and many with low frequency.
 Figure 2.
-The Venn diagram reveals that a core set of 5672 headers are common across Chrome, Firefox, and Brave browsers across two years.
-Unique headers are present in each browser, with $\text{Chrome}_{23}$ having the highest number of unique headers, which may reflect browser-specific features or experimental headers.
-Firefox had a higher header similarity with Chrome than with Brave.
-Figure 2 represents a Venn diagram and provides a direct comparison of response headers between different browser versions, highlighting data and feature similarity.
+Venn diagram of HTTP response headers in the examined datasets, revealing a core set of 5672 headers that are common across Chrome, Firefox, and Brave.
+Unique headers are present in each browser, although $\text{Chrome}_{23}$ has the highest number of unique headers, which may reflect browser-specific features or experimental headers.
+Figure 2 depicts a Venn diagram that compares the response headers present in the examined datasets, highlighting data and feature similarity between the different browsers.
 The intersection numbers, such as the 5672 headers common to all browsers, suggest a significant overlap in features across different platforms.
-However, the presence of headers unique to each browser or shared by only two browsers (e.g., 1427 unique to $\text{Chrome}_{22}$ and 154 unique to $\text{Brave}_{22}$) indicates that while there is a core set of features, there is also a notable amount of variation which can be critical for web tracker detection.
-The Empirical Cumulative Distribution Function plot for $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$ in Figure 3 row (A) show that the Content-Length values for non-trackers are generally higher than those for trackers, which suggests that non-trackers tend to send larger responses.
-This is consistent across both browser versions, indicating that while browsers evolve, certain characteristics like the smaller size of tracker responses remain consistent.
-The second row (B) plots the feature similarity for the X-XSS-Protection header, showing that header values can be very similar between the datasets, i.e., the training and test sets.
-It further supports our observation that the datasets are similar and that trackers and non-trackers have differences across the usage of HTTP header fields and values.
-Although we present only few selected examples here, they should highlight the general challenges that come with response messages datasets.
+However, the presence of headers unique to each browser or shared by only two browsers indicates that while there is a core set of shared features, there is also a notable amount of variation that may be critical for web tracker detection.
+For instance, the Empirical Cumulative Distribution Function plots for $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$ presented in Figure 3 row (A) show that the Content-Length values of non-trackers are generally higher than those of trackers, which suggests that non-trackers tend to send larger responses.
+This is consistent across both datasets, indicating that while browsers evolve, certain characteristics, like the smaller size of tracker responses, remain consistent.
+Conversely, row (B) of Figure 3 plots the feature similarity for the X-XSS-Protection header, showing that header values can be very similar between the datasets, i.e., the training and test sets.
+This further supports our observation that most of the examined datasets are similar and that trackers and non-trackers differ in their use of HTTP header fields and values.
+Although we have presented only a few selected examples here, they highlight the general challenges that come with response message datasets.
 Figure 3.
 Row (A) presents the ECDFs for both Chrome datasets for the Content-Length header values.
-We set a cut-off value at 10 000 to highlight the core observation and more than half of the responses have a value are below this threshold – $\tilde{\mu}_{T22}=62$, $\tilde{\mu}_{NT22}=8068$ and $\tilde{\mu}_{T23}=45$, $\tilde{\mu}_{NT23}=8021.5$.
-Row (B) shows how similar the values across Chrome, Firefox, and Brave are for the X-XSS-Protection header.
-Additional characteristics of the datasets include the high amount of missing values per header.
-One explanation might be that servers do not need to use every header field that exists, and it is common practice to use developer headers that are not part of the standard, i.e.
-they are not registered in the Request for Comments 4229111RFC 4229, https://www.ietf.org/rfc/rfc4229.txt, accessed: November 30, 2023.
-There are multiple headers that contain mostly unique values, e.g., Date.
-Lastly, according to the used filter lists, most HTTP response messages are not related to any tracking activities.
-This entails a class imbalance to be handled carefully to derive meaningful results from the final models.
+We set a cut-off value at 10,000 to highlight the core observation and more than half of the responses have a value below this threshold – $\tilde{\mu}_{T22}=62$, $\tilde{\mu}_{NT22}=8068$ and $\tilde{\mu}_{T23}=45$, $\tilde{\mu}_{NT23}=8021.5$.
+Row (B) shows how similar the values across $\text{Chrome}_{22}$, $\text{Firefox}_{22}$, and $\text{Brave}_{22}$ are for the X-XSS-Protection header.
+Additional characteristics of the datasets include a high amount of missing header values, resulting from a large number of available HTTP headers, of which any given response only uses a small number.
+This is exacerbated by the common practice of using custom headers that are not part of the HTTP standard, i.e., are not registered in RFC 4229555https://www.ietf.org/rfc/rfc4229.txt.
+This results in the presence of 8058 custom headers in the $\text{Chrome}_{22}$ dataset, contrasted by a comparatively small $\tilde{\mu}$ of unique headers per response as shown in Table 1.
+The examined datasets are formatted as a matrix, with rows representing individual HTTP requests or responses and each unique HTTP header denoted in a separate column.
+This, combined with the observed large total number of unique headers and small number of headers per response, results in a sparse matrix where most elements are empty.
 5. Methodology
-Based on our findings from the data exploration in Section 4, we propose our methodology to assess the effectiveness of classifiers based on response headers.
+Based on the findings of our data exploration presented in Section 4, we propose a comprehensive methodology to assess the effectiveness of classifiers based on response headers.
 5.1. Problem Formulation
 We consider the problem of binary classification, where we aim to predict the presence of web trackers based on response headers.
-The headers, represented by the feature vector $x\in\mathbb{R}^{d}$, are nominal and can be both discrete and continuous variables.
-These variables are derived from HTTP headers of four different browser types, denoted as the set $B=\{\text{Chrome},\text{Firefox},\text{Brave}\}$.
-For this study, we specifically consider versions $\text{Chrome}_{22},\text{Chrome}_{23},\allowbreak\text{Firefox}_{22},\text{Brave}_{22}$.
-Given a training set $\mathcal{D}_{\text{train}}=\{(x_{1}y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}$ is a feature vector of the $i$-th HTTP header and $y_{i}\in\{0,1\}$ represents the associated class label (with ’1’ indicating the presence of a web tracker), our goal is to train a set of classifiers $f_{j}:\mathbb{R}^{d}\rightarrow\{0,1\}$ for $j\in M$ that minimizes the empirical risk:
+The headers, represented by the feature vector $x\in\mathbb{R}^{d}$, are nominal and can be discrete and continuous variables.
+These variables are derived from HTTP headers of four different browser types, denoted as the previously defined set $B$.
+For this study, we specifically consider the datasets Chrome22, Chrome23, Firefox22, and Brave22.
+Given a training set $\mathcal{D}_{\text{train}}=\{(x_{1}y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}$ is a feature vector of the $i$-th HTTP header and $y_{i}\in\{0,1\}$ represents the associated class label, our goal is to train a set of classifiers $f_{j}:\mathbb{R}^{d}\rightarrow\{0,1\}$ for $j\in M$ that minimizes the empirical risk:
 (1) $R(f_{i})=\frac{1}{N}\sum_{i=1}^{N}L(y_{i},f_{i}(x_{i})),\quad j\in\{1,\ldots,M\},$
 where $L(y_{i},f_{i}(x_{i}))$ is the loss incurred by predicting $f_{i}(x_{i})$ when the true label is $y_{i}$.
-The specific form of the loss function $L(\cdot)$ will be dependent on the particular classifier $f_{j}$ used.
-We consider datasets represented by $D_{s}$, where $s$ can pertain to training, testing, validation, or calibration subsets of the data. We observe several challenges from our data exploration:
-  1. (1)
-    High Cardinality: The feature vectors $x_{i}$ in our datasets $D_{s}$ can exhibit high cardinality for certain features. This problem can pose challenges in terms of model training and encoding.
-  2. (2)
-    Sparseness: The feature vectors $x_{i}$ in our datasets $D_{s}$ are represented as sparse vectors, with each vector defined as:
+The specific form of the loss function $L(\cdot)$ will be dependent on the classifier $f_{j}$ used.
+We consider datasets represented by $D_{s}$, where $s$ can pertain to training, testing, or calibration subsets of the data.
+We observe several challenges (CH) and key characteristics from our data exploration, which answer our RQ1:
+  • (CH1) High Cardinality: The feature vectors $x_{i}$ in our datasets $D_{s}$ may exhibit high cardinality for certain features.
+This problem may pose challenges in terms of model training and encoding.
+  • (CH2) Sparseness: As discussed above, the feature vectors $x_{i}$ in our datasets $D_{s}$ are represented as sparse vectors, with each vector defined as:
     $x_{i}=\{v_{1},v_{2},\ldots,v_{d}\}$
-    Here, a significant proportion of the entries $v_{j}$ within $x_{i}$ are either zero or represent missing data, leading to the overall sparseness of $D_{s}$.
-Such sparseness can introduce computational complexities and may challenge the performance of certain machine learning algorithms not optimized for sparse data.
-  3. (3)
-    Class Imbalance: For datasets collected from a given browser $B$ and represented as $D_{s}$, the class distribution can exhibit imbalance.
+    A significant proportion of the entries $v_{j}$ within $x_{i}$ are either zero or represent missing data, resulting in an overall sparseness of $D_{s}$.
+Such sparseness may introduce computational complexities and challenge the performance of certain ML algorithms that are not optimized for sparse data.
+  • (CH3) Class Imbalance: For datasets collected from a given browser $B$ and represented as $D_{s}$, the class distribution may be imbalanced.
 The majority class is denoted by $\alpha_{B}$, and the minority class is represented by $\beta_{B}$, satisfying $\alpha_{B}+\beta_{B}=1$.
-The degree of this imbalance can vary based on the browser as shown in Table 1, with some showing a more pronounced disparity between $\alpha_{B}\text{ and }\beta_{B}$.
+The degree of this imbalance may vary based on the browser, as shown in Table 1, with some datasets exhibiting a more pronounced disparity between $\alpha_{B}\text{ and }\beta_{B}$.
 Such imbalance can introduce bias to models, favoring the majority class and potentially undermining the accurate prediction of the minority class.
-  4. (4)
-    High Dimensionality: With a high dimensionality $d$ and a high number of observations $n$ per dataset, the classification is susceptible to the Curse of Dimensionality, potentially leading to overfitting and necessitating intricate dimensionality reduction techniques.
-We have to apply a dimensionality reduction process that achieves a reduced feature space $X_{i}^{\prime}\subseteq X$, entailing a reasonable training time.
-  5. (5)
-    Data Similarity: The datasets $\mathcal{D}_{s}$ collected across different browsers can exhibit pronounced similarities in their feature distributions and values.
+  • (CH4) High Dimensionality: With a high dimensionality $d$ and a high number of observations $n$ per dataset, the classification is susceptible to the Curse of Dimensionality, potentially leading to overfitting and necessitating intricate dimensionality reduction techniques.
+Therefore, applying a process that achieves a reduced feature space $X_{i}^{\prime}\subseteq X$ is necessary, entailing a reasonable training time.
+  • (CH5) Data Similarity: The datasets collected across browsers can exhibit pronounced similarities in their feature distributions and values.
 Given the high dimensionality of our feature space, exhaustive pairwise comparisons can be computationally expensive and time-consuming, highlighting the challenge of accurately gauging data similarities in such a complex space.
+Figure 4. The parameterized pipeline automatically processes the collected datasets and trains the classifiers.
 5.2. Machine Learning Pipeline
-Our machine learning pipeline, which we present in the following, addresses the previously stated problems for the classification of web trackers.
-Many different approaches have been applied in order to uncover web trackers.
-Another closely related field to web privacy is web security research.
-Although they seek to detect different entities, both have the exploitation of HTTP data in common.
-Typically, a dataset goes through extensive data processing and feature engineering to find useful features, which includes a thorough analysis of the values within each feature.
-Despite this common conception, Laughter et al.
-(Laughter et al., 2021) were interested if features that merely represented the presence of an HTTP header in a binary fashion were usable in a classifier to detect malicious HTTP requests based on HTTP headers and URLs.
-In the end, some of their classifiers performed well with accuracies above 90%.
-Similar results were observed in (Bhagavatula et al., 2014) and partly by (Dudykevych and Nechypor, 2016).
-This method is interesting on two levels: (i) it cuts down the time of feature engineering, which is fairly complex in this use case as HTTP headers are mainly categorical data, and (ii) it should help when the models are deployed and used in production, e.g.
-if HTTP data is received, it should be evaluated in real time and checking the presence of headers is faster than elaborate processing before evaluation.
-Therefore, the binary representation of HTTP headers as a feature will be tested.
-First and foremost, we train our classifiers based on the $\text{Chrome}_{22}$ dataset and use all other datasets as test sets, which is in accordance with RQ3.
-Generally, each step in our pipeline focuses on reducing the dimensions of the datasets, as most headers are rarely used, thus reducing potential overfitting.
-This approach further reduces the sparseness and potential features with high cardinality.
-The latter challenge is solved by employing the binary-encoding method as it flattens the respective feature to two possible values.
-The sparseness and high dimensionality are related to each other, in a sense that most headers and their corresponding values are rare or even unique.
-We apply multiple methods to reduce the dimensionality: (i) a low variance filter to remove headers with only one value, (ii) the removal of headers with a missing value ratio of one, (iii) headers related to only one label are dropped, and (iv) Approximate string matching, sometimes referred to as Fuzzy Matching to combine header fields with a similarity of at least 80% based on a weighted ratio of multiple distance metrics, e.g., Damerau Levenshtein or Hamming.
-All of this reduces the sparseness further and binary-encoding simplifies the sparse matrix.
-Although a multitude of different resampling techniques exists to combat class imbalance, the question of whether this step is actually necessary remains.
+This subsection outlines our ML pipeline that addresses the challenges in classifying web trackers.
+Related work uses various approaches to detect web trackers, a task closely related to web security research.
+Both research fields exploit HTTP data, typically requiring extensive data processing and feature engineering to identify useful features.
+However, Laughter et al.
+(Laughter et al., 2021) explored whether binary features that represent the mere presence of HTTP headers could effectively detect malicious HTTP requests.
+Their classifiers achieved accuracy scores above 90%.
+Similar findings were reported by (Bhagavatula et al., 2014) and (Dudykevych and Nechypor, 2016).
+The binarization of HTTP headers is appealing for two reasons: (i) it simplifies feature engineering, which is complex for categorical header data, and (ii) it enhances model deployment for tracker detection, as checking header presence is faster than detailed processing.
+Therefore, we apply and evaluate this method in our analysis.
+We train our classifiers on the $\text{Chrome}_{22}$ dataset, using other datasets for testing as per RQ3.
+Therefore, only the training set is used to filter out headers, and the results of all steps in the data processing phase in Figure 4 are afterwards applied to all testing sets.
+Each step in our pipeline aims to reduce dataset dimensions, addressing overfitting and sparsity issues.
+To handle sparsity and high dimensionality and simplify the data matrix, we use several methods: (i) applying a low variance filter to remove headers with a single value, (ii) eliminating rare headers with a missing-value ratio approaching one, (iii) dropping headers associated with only one label as they might be correlated to the label, and (iv) employing Approximate string matching (Fuzzy Matching) to merge similar headers based on a weighted ratio of distance metrics (e.g., Damerau Levenshtein and Hamming).
+Fuzzy matching is especially useful due to the presence of hundreds of similar headers within our data.
+Our initial data analysis discussed in Section 4 identifies many examples where headers with similar purpose were written or formatted differently, e.g., content-lenght, cteonnt-length, or ntcoent-length, instead of content-length.
+Therefore, comparing the header names is insufficient, so we automatically assess the value similarity to reduce potential mismatches.
+Although many different resampling techniques exist to combat class imbalance, such as over- and undersampling (Hernandez et al., 2013), generative adversarial networks (Engelmann and Lessmann, 2021), or variational autoencoders (Fajardo et al., 2021), the question of whether this step is necessary remains.
 Some disadvantages of such methods are, for example, the removal of observations, the replication of existing observations, or the creation of synthetic samples.
-Selecting and applying a sampling method can lead to overfitting and overoptimism as well (Santos et al., 2018), notwithstanding the question of whether it is even feasible to train the models on a distribution that was not observed in real-world data.
-Our pipeline preserves the observed distribution as the imbalance is mild (goo, 2023) in the $\text{Chrome}_{\textbf{22}}$ dataset, i.e., a sufficient number of samples for each class exists, which allows proper training to differentiate between classes.
-Furthermore, we compute metrics that consider or are not affected by the imbalance of the datasets.
-The same applies for most of the selected ensemble models.
-Fuzzy matching is especially useful due to the presence of hundreds of similar headers within our data.
-During our initial data analysis deiscussed in Section 4, we identified many examples where headers with similar purpose were written or formatted differently, e.g.
-content-lenght, cteonnt-length, or ntcoent-length, instead of content-length.
-Accordingly, simply comparing the header names is not enough, which is why we automatically assess the value similarity to reduce potential mismatches.
-Lastly, we apply Isotonic Regression instead of Platt Scaling to calibrate the probabilities outputted by our models, following the results from (Niculescu-Mizil and Caruana, 2005).
-This is done by fitting an Isotonic Regression model to the models’ predictions and the true labels, and using this to adjust the predicted probabilities.
-The calibration function for model $i$ is denoted as $C_{i}(f_{i}(x))$.
-We summarise our pipeline in Figure 4, including the data collection.
+Selecting and applying a sampling method can lead to overfitting and over-optimism as well (Santos et al., 2018), notwithstanding the question of whether it is even feasible to train the models on a distribution that was not observed in real-world data.
+Addressing CH3, our pipeline preserves the observed distribution as the imbalance is mild  [^6] https://developers.google.com/machine-learning/data-prep/construct/sampling-splitting/imbalanced-data in the $\text{Chrome}_{22}$ dataset, i.e., a sufficient number of samples exists for each class, which allows proper training to differentiate between classes.
+We also compute metrics that account for or are unaffected by imbalance, and the selected ensemble models are inherently robust to this issue.
+Lastly, following the findings of Niculescu-Mizil and Caruana (Niculescu-Mizil and Caruana, 2005), we apply Isotonic Regression instead of Platt Scaling to calibrate the probabilities calculated by our models.
+This is done by fitting an Isotonic Regression model to the models’ predictions as well as the true labels to adjust the predicted probabilities.
+The calibration function for each model $i$ is denoted as $C_{i}(f_{i}(x))$.
+We summarize our pipeline in Figure 4, including the data collection.
 Note that we did not apply any data imputation to keep the final datasets as close to the original collected datasets as possible.
-A validation set for hyperparameter tuning was not created because we want to assess the effect of calibration on its own.
-Moreover, tuning requires substantially more computational power, e.g.
-by using Grid Search.
-Figure 4. The parameterized pipeline automatically processes the collected datasets and trains the classifiers.
+In addition, due to the application of binarization as a final step, we do not need to impute values for missing data as we are solely interested in the presence of headers and not their values.
+We do not generate a validation set for hyperparameter tuning because we want to assess the effect of calibration in isolation.
+Moreover, tuning requires more computational power.
+However, this might lead to suboptimal configurations of our classifiers, thus a reduced performance.
 5.3. Classification Models
-As discussed in Section 2.1, prior studies (Cozza et al., 2020; Li et al., 2015; Bhagavatula et al., 2014; Gugelmann et al., 2015) often used similar classifiers despite the broad range of available models.
+As discussed in Section 2.1, related studies (Cozza et al., 2020; Li et al., 2015; Bhagavatula et al., 2014; Gugelmann et al., 2015) often used similar classifiers despite the broad range of available models.
 We build upon these established selections, refining our choice of models based on three conditions: (i) the models should be trainable in a reasonable amount of time, (ii) they should be applicable to this specific use case, and (iii) they should preferably be ensemble methods, although we include a handful of non-ensemble models that were used in past research to allow for direct comparisons between our results and those of related studies.
-We prefer ensemble methods as they usually perform better than their counterparts (Dietterich, 2000), leading to their popularity over the past few years.
-Nevertheless, to our best knowledge, they have not been widely used in past web privacy research, except for Random Forests (RF).
+We prefer ensemble methods, as they usually perform better than their counterparts (Dietterich, 2000), leading to their popularity over the past few years.
+Nevertheless, to the best of our knowledge, they have not been widely used in past web privacy research, with the exception of Random Forests.
 In Table 2, we provide an overview of our selected models and their use in related work.
@@
 Table 2.
-Overview of the selected machine learning models and the number of times they were used in past research from Section 2.1.
-One paper did not exactly specified which ensemble model was used, thus, the most likely model was marked with Unspecified (U).
-Papers from web security are underlined.
-Other supervised learning methods, e.g., Multi-layer Perceptrons (Cozza et al., 2020; Guarino.
+Overview of the selected ML models and their use in related research discussed in Section 2.1.
+One paper did not specify exactly which ensemble model was used.
+The most likely model is marked as Unspecified (U).
+Papers from the field of web security are underlined.
+Other supervised learning methods, such as Multi-layer Perceptrons (Cozza et al., 2020; Guarino.
 et al., 2020), KNN, and especially SVM, could have been used.
@@
 et al., 2020).
-However, both suffer in terms of time and computational complexity, e.g.
-SVM has a train time complexity of $O(n^{2})$ where $n$ is the number of observations (Steinwart and Christmann, 2008).
-In addition, deep-learning models are not considered as the focus is on supervised classifiers and simple rather than complex models.
-A graph-based approach, used by several papers (Kalavri et al., 2016; Castell-Uroz et al., 2021) will not be pursued because it does not fit well with the targeted approach, and direct connections between individual HTTP responses will not be analyzed, e.g.
-through grouping based on their URL or building a graph-based on their order.
+However, both suffer in terms of time and computational complexity.
+SVM, for instance, has a training time complexity of $O(n^{2})$, where $n$ is the number of observations (Steinwart and Christmann, 2008).
+In addition, deep-learning models are not considered, as our research focuses on simple, lightweight models.
+We also do not pursue a graph-based approach as used by several papers (Kalavri et al., 2016; Castell-Uroz et al., 2021) because we do not examine connections between responses, e.g., through grouping based on their URL or building a graph based on their order.
 5.4. Baseline Models
 We compare our classifiers against three baseline models and multiple metrics that we define in Section 5.6.
-We select one tree classifier, the Decision Tree (DT), that is also the building block of six selected classification models.
-Two simplistic and relatively robust baselines, namely Naïve Bayes (NB) and Logistic Regression (LR).
-Additionally, it is good practice to compare against a baseline approach that differs from the presented classifiers while using the same dataset.
+We select one tree classifier, the Decision Tree (DT), which is the foundation of six selected classification models.
+Our other two baselines are Gaussian Naïve Bayes (GNB) and Logistic Regression (LR), two simplistic and relatively robust models.
+It is good practice to compare against a baseline approach that differs from the presented classifiers while using the same dataset.
 Raschke et al.
-(Raschke et al., 2023) presented a graph-based approach named T.EX-Graph where edge and node attributes as well as centrality metrics served as features.
-Furthermore, the authors included information from the responses but did not pursue classifiers that are solely based on response headers.
-This comparability should help in assessing the performance gains and the relevance of response header values for automated web tracker detection.
-However, the authors applied Synthetic Minority Oversampling to balance their dataset.
-We focus on their results using the FQDN dataset, where hosts are modeled as fully qualified domain names (FQDN) because the tracker distribution is similar to ours with a NT to T ratio of 68:32.
+(Raschke et al., 2023) presented a graph-based approach named T.EX-Graph where edge and node attributes and centrality metrics served as features.
+Furthermore, the authors included information from the responses but did not pursue classifiers based solely on response headers.
+This comparability should help assess the performance gains and the relevance of response header values for automated web tracker detection, although the authors applied Synthetic Minority Oversampling to balance their dataset.
+We focus on their results stemming from the FQDN dataset, where hosts are modeled as fully qualified domain names (FQDN) because the tracker distribution is similar to that of our datasets with a non-tracker to tracker ratio of 68:32.
+To verify the results and calculate additional metrics not used in the original study, we replicated their experiment.
+5.5. Models using HTTP Request Headers
+Lastly, we compare our proposed approach to HTTP request header-only models.
+Our reasoning for this comparison is twofold: (i) showing which additional information can be gained from HTTP response headers, thus the usefulness of our approach and (ii) whether the applied technique of binarization is helpful for models using HTTP request headers.
+To enable a direct comparison between request- and response-based classifiers, we apply the pipeline presented in Section 5.2 to the request headers present in our datasets and compare the results for each request-response pair.
 5.6. Evaluation Metrics
-We measure the performance of our classifiers according to nine different metrics.
-Several metrics are common in the web tracking literature, e.g., accuracy, precision, or F1-Score, while others were selected due to their applicability to our classification problem and to understand the relationship with the underlying data in more detail.
-Additional metrics are reported in the Appendix, e.g., the confusion matrix because its interpretation is already captured by other metrics more concisely.
-Moreover, we calculate confidence intervals (CI) for these metrics using bootstrap sampling with 599 samples (Wilcox, 2010).
+We measure the performance of our classifiers according to nine metrics.
+Several of these metrics are common in the web tracking literature, while others are selected due to their applicability to our classification problem and to understand the relationship with the underlying data in more detail.
+We report on additional metrics, such as the confusion matrix, in the appendix.
+Moreover, we calculate confidence intervals (CI) for these metrics using bootstrap sampling with 599 samples (Wilcox, 2001).
 Metric Past Research
@@
 MCC (McGahagan et al., 2019)
-AUCPRC Not used
+AUPRC Not used
 Log-Loss Not used
-Table 3. Overview of performance metrics and the number of times they were used in past research from Section 2.1.
-The first metrics that are commonly used in web privacy research are Accuracy, Precision, Recall, False Positives (FP), True Positives (TP), False Negatives (FN), and True Negatives (TN).
-Two accuracy scores can be calculated, one from the perspective of trackers, the positive label, and one from non-trackers, the negative label.
-However, accuracy is not particularly useful with imbalanced datasets because a model could always predict the majority class and achieve accuracy with the same value as the distribution of said class.
-We calculated additional, more reliable metrics to quantify the performance of a classifier with an imbalanced dataset as follows:
+Table 3. Overview of applied performance metrics and their use in related research discussed in Section 2.1.
+The first metrics that are commonly used in web privacy research are Accuracy, Precision, Recall, False Positives (FP), True Positives (TP), False Negatives (FN), True Negatives (TN), and F1-Score.
+Two accuracy scores can be calculated: one from the perspective of trackers, the positive label, and one from non-trackers, the negative label.
+However, accuracy is not useful with imbalanced datasets because a model could always predict the majority class and achieve accuracy with the same value as the distribution of said class.
+For this reason, we calculate the following additional metrics to quantify the performance of a classifier with an imbalanced dataset:
 Balanced Accuracy:
 Broderson et al.
-(Brodersen et al., 2010) introduced the Balanced Accuracy (BACC) metric for imbalanced datasets in 2010 as an alternative to average accuracy scores, which are, according to the authors, problematic for confidence intervals (CI) and imbalanced datasets.
-The metric comprises recall and TNR and is defined as:
+(Brodersen et al., 2010) introduced the Balanced Accuracy (BACC) metric for imbalanced datasets in 2010 as an alternative to average accuracy scores, which are, according to the authors, problematic for CIs and imbalanced datasets.
+The metric comprises recall and true negative rate (TNR) and is defined as:
 (2) $BalancedAccuracy=\frac{1}{2}\left(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\right)$
-F1-Score:
-Another metric that is commonly used for classification models is the F1-Score which expresses a model’s performance in one value, despite the ambivalent nature of precision and recall.
-Compared with BACC, it emphasizes positives more, whereas BACC takes negatives more into account, which can result in two very different results.
-The F1-Score is mathematically defined as the harmonic mean between recall and precision:
-(3) $F_{1}\text{-}Score=\frac{2\cdot Precision\cdot Recall}{Precision+Recall}$
 Log-loss Score:
-This metric, sometimes referred to as cross-entropy-loss, is particularly useful for classification problems as it focuses on the predictive performance of a model.
-The function is applied to each observation individually, and then the average over all elements is calculated.
-It describes how large the error between a prediction probability and the true label is, and a value closer to zero represents a very good score.
-However, this metric can be influenced by the imbalance in a dataset, i.e.
-the baseline log-loss gets smaller with an increase in imbalance, and the log-loss metric should always be compared to its baseline value, e.g.
-the baseline log-loss for a dataset with a 70:30 ratio equals 0.611, and 0.325 for datasets with a 90:10 ratio (tds, 2023).
-All in all, the model’s log-loss should be lower than its baseline.
-The math formula is defined as:
+Also referred to as cross-entropy-loss, the log-loss score focuses on the predictive performance of a model.
+It describes how large the error between a prediction probability and the true label is, where a value closer to zero represents a better score.
+However, this metric can be influenced by an imbalance, i.e., the baseline log loss gets smaller with an increase in imbalance.
+Therefore, it is essential to compare the metric to its baseline value.
+For instance, the baseline for our $\text{Chrome}_{22}$ dataset with a 70:30 ratio equals 0.611  [^7] https://towardsdatascience.com/intuition-behind-log-loss-score-4e0c9979680a.
+The model’s log-loss should be lower than its baseline.
+The log-loss score is defined as:
 (3) $L_{log}(y,p)=-[y\cdot log(p)+(1-y)\cdot log(1-p)]$
-where $y$ stands for the label and $p$ for probability.
+where $y$ is the label class and $p$ is the probability.
 Matthews-Correlation-Coefficient:
-One metric is often not enough to fully comprehend the performance of a model and if it is even suited for binary classification problems.
-Matthews-Correlation-Coefficient (MCC) is becoming more popular in recent years, although it was introduced in 1975 (Matthews, 1975) – it is defined as:
+The Matthews-Correlation-Coefficient (MCC) (Matthews, 1975), a metric that has seen increasing popularity in recent years, is defined as:
 (4) $MCC=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)\cdot(TP+FN)\cdot(TN+FP)\cdot(TN+FN)}}$
-However, it is rarely used in web privacy research compared to other fields like web security research.
-MCC ranges from -1 to +1 and can be understood as a correlation coefficient between the true and predicted label.
-One particular advantage is the inclusion of values from the confusion matrix, i.e.
-the MCC will only achieve a high value if both class labels were predicted with high accuracy – which in turn presents the MCC as a meaningful metric for imbalanced datasets (Boughorbel et al., 2017).
-This was further substantiated by Chicco et al.
-(Chicco and Jurman, 2020) in 2020, and the authors outlined how accuracy and F1-Score were not reliable measures for imbalanced datasets.
-ROC-AUC:
-Our next metric is the Receiver Operating Characteristics (ROC) curve, complemented by the Area Under the curve (AUC) value.
-The former describes a classifier’s performance at various thresholds by plotting the TPR on the ordinate and FPR on the abscissa.
-The latter metric expresses how well the classifier is able to separate between both class labels.
+MCC ranges from -1 to +1 and can be understood as a correlation coefficient between the true and predicted labels.
+The MCC will only achieve a high value if both class labels are predicted with high accuracy, positioning the MCC as a meaningful metric for imbalanced datasets (Boughorbel et al., 2017).
+This was further substantiated in 2020 by Chicco et al.
+(Chicco and Jurman, 2020), who outlined how accuracy and F1-Score are not reliable measures for imbalanced datasets.
 AUPRC:
 The last metric in our study is the Area Under the Precision-Recall Curve (AUPRC).
-This metric is particularly relevant in the context of a supervised binary classification problem, especially when dealing with imbalanced datasets.
-The Precision-Recall curve is a plot that illustrates the trade-off between the precision (positive predictive value) and the recall (true positive rate) for a classifier at various thresholds.
-Precision is defined as the ratio of true positives to the sum of true and false positives, while recall is the ratio of true positives to the sum of true positives and false negatives.
+This metric is relevant when dealing with imbalanced datasets.
+The Precision-Recall Curve is a plot that illustrates the trade-off between precision and recall for a classifier at various thresholds.
 The AUPRC is an aggregated measure of this curve, representing the classifier’s ability to correctly identify positive instances across different thresholds.
-In imbalanced settings, where one class is significantly underrepresented, the AUPRC is a more informative and reliable metric than ROC-AUC, as it focuses on the performance of the classifier on the minority class, which is often of greater interest (Saito and Rehmsmeier, 2015).
+In imbalanced settings, where one class is significantly underrepresented, AUPRC is a more informative and reliable metric than ROC-AUC, as it focuses on the performance of the classifier on the minority class, which is of interest for us (Saito and Rehmsmeier, 2015).
 6. Experiments
+Figure 5.
+t-SNE plots using a representative sample across all browsers to gauge data similarity.
+Separate clusters exist for trackers and non-trackers, but there is significant overlap between both classes.
+The results are separated by class to highlight differences between trackers and non-trackers.
+In this section, we evaluate the effectiveness of our proposed methodology in detecting trackers.
+We begin by presenting our classifier results based on $\text{Chrome}_{22}$ data and follow up by testing them with data from $\text{Firefox}_{22}$ and $\text{Brave}_{22}$.
+Furthermore, we assess the longitudinal performance with the $\text{Chrome}_{23}$ dataset that was collected almost a year after $\text{Chrome}_{22}$.
+Note that this section only summarizes our results; performance metrics across all classifiers are reported in the Appendix 9.1 and 9.2.
+We test our approach on a MacBook Pro (14-inch, 2021) with a 10-core CPU, a 16-core GPU, 1TB of SSD storage, and 32GB of RAM.
+We address CH5 by applying t-SNE to our datasets as shown in Figure 5.
+Non-trackers form high-density clusters, indicating that they share similar headers.
+$\text{Brave}_{22}$ exhibits more smaller groups at the edges that are separated from the others.
+Trackers only form one large cluster, showing that commonalities exist even though they are more dispersed overall.
+This is potentially due to the higher number of custom headers present in tracker responses.
+Additionally, overlaying the t-SNE plots for trackers and non-trackers reveals a substantial overlap between the two classes.
 Model Accuracy Log-Loss ROC-AUC AUPRC BACC F1-Score Precision Recall MCC
-LR 0.892 0.252 0.86 0.902 0.86 0.812 0.847 0.78 0.738
-[0.891;0.894] [0.25;0.255] [0.858;0.862] [0.9;0.904] [0.858;0.862] [0.81;0.815] [0.843;0.85] [0.777;0.784] [0.735;0.742]
-GNB 0.853 2.47 0.832 0.792 0.832 0.759 0.74 0.779 0.654
-[0.851;0.854] [2.436;2.514] [0.829;0.833] [0.788;0.797] [0.829;0.833] [0.756;0.762] [0.737;0.745] [0.776;0.782] [0.65;0.657]
-BNB 0.851 0.617 0.833 0.849 0.833 0.759 0.732 0.789 0.652
-[0.849;0.852] [0.608;0.627] [0.83;0.835] [0.847;0.852] [0.83;0.835] [0.757;0.763] [0.729;0.736] [0.784;0.793] [0.648;0.657]
-DT 0.956 0.546 0.947 0.943 0.947 0.926 0.928 0.925 0.895
-[0.955;0.957] [0.531;0.566] [0.946;0.948] [0.94;0.945] [0.946;0.948] [0.925;0.928] [0.926;0.93] [0.922;0.927] [0.893;0.897]
-RF 0.96 0.161 0.951 0.978 0.951 0.932 0.935 0.929 0.903
-[0.959;0.96] [0.155;0.168] [0.949;0.952] [0.977;0.979] [0.949;0.952] [0.93;0.934] [0.933;0.937] [0.926;0.931] [0.901;0.906]
-ET 0.964 0.166 0.953 0.98 0.953 0.938 0.952 0.925 0.913
-[0.963;0.965] [0.159;0.174] [0.952;0.954] [0.978;0.98] [0.952;0.954] [0.937;0.94] [0.95;0.954] [0.923;0.927] [0.911;0.915]
-AdaBoost 0.888 0.574 0.856 0.894 0.856 0.806 0.838 0.776 0.729
-[0.887;0.89] [0.574;0.575] [0.854;0.858] [0.892;0.897] [0.854;0.858] [0.803;0.809] [0.834;0.841] [0.772;0.78] [0.725;0.733]
-GBM 0.895 0.244 0.867 0.916 0.867 0.82 0.845 0.795 0.747
-[0.894;0.897] [0.242;0.246] [0.865;0.869] [0.914;0.918] [0.865;0.869] [0.817;0.822] [0.842;0.848] [0.792;0.799] [0.743;0.75]
-LGBM 0.931 0.174 0.909 0.957 0.909 0.88 0.908 0.854 0.833
-[0.93;0.932] [0.172;0.175] [0.907;0.91] [0.956;0.958] [0.907;0.91] [0.878;0.883] [0.906;0.911] [0.851;0.857] [0.83;0.836]
-HistGB 0.929 0.175 0.907 0.956 0.907 0.877 0.902 0.853 0.827
-[0.927;0.93] [0.173;0.176] [0.905;0.909] [0.955;0.957] [0.905;0.909] [0.875;0.879] [0.9;0.905] [0.85;0.856] [0.825;0.831]
-XGBoost 0.928 0.172 0.912 0.956 0.912 0.878 0.885 0.872 0.827
-[0.927;0.93] [0.17;0.174] [0.91;0.914] [0.955;0.957] [0.91;0.914] [0.877;0.881] [0.882;0.889] [0.869;0.875] [0.825;0.831]
-XGBoost 0.883 - - - - 0.867 0.867 0.867 -
-RF 0.881 - - - - 0.866 0.862 0.871 -
-GBM 0.858 - - - - 0.844 0.836 0.856 -
-SVC 0.848 - - - - 0.828 0.819 0.852 -
-AdaBoost 0.840 - - - - 0.824 0.816 0.836 -
-DT 0.834 - - - - 0.814 0.810 0.819 -
-KNN 0.833 - - - - 0.814 0.809 0.820 -
-LR 0.809 - - - - 0.796 0.789 0.822 -
+LR 0.874 0.321 0.829 0.858 0.829 0.773 0.836 0.718 0.69
+[0.872;0.875] [0.318;0.324] [0.827;0.831] [0.856;0.862] [0.827;0.831] [0.769;0.776] [0.831;0.839] [0.714;0.723] [0.686;0.693]
+GNB 0.835 3.232 0.803 0.688 0.803 0.723 0.723 0.723 0.605
+[0.832;0.836] [3.193;3.28] [0.8;0.805] [0.683;0.693] [0.8;0.805] [0.72;0.727] [0.719;0.727] [0.719;0.727] [0.6;0.61]
+DT 0.958 0.324 0.945 0.965 0.945 0.929 0.945 0.913 0.9
+[0.957;0.959] [0.311;0.34] [0.944;0.946] [0.963;0.966] [0.944;0.946] [0.928;0.93] [0.943;0.947] [0.911;0.915] [0.898;0.902]
+RF 0.959 0.162 0.946 0.98 0.946 0.93 0.947 0.913 0.901
+[0.958;0.96] [0.156;0.17] [0.944;0.947] [0.979;0.98] [0.944;0.947] [0.928;0.931] [0.949;0.949] [0.911;0.915] [0.899;0.903]
+ET 0.96 0.195 0.946 0.977 0.946 0.931 0.951 0.912 0.903
+[0.959;0.961] [0.186;0.204] [0.945;0.947] [0.976;0.978] [0.945;0.947] [0.93;0.933] [0.949;0.953] [0.91;0.914] [0.901;0.906]
+AdaBoost 0.858 0.673 0.809 0.846 0.809 0.742 0.809 0.686 0.65
+[0.856;0.859] [0.673;0.673] [0.806;0.811] [0.844;0.849] [0.806;0.811] [0.739;0.746] [0.805;0.813] [0.682;0.69] [0.646;0.653]
+GBM 0.892 0.292 0.85 0.89 0.85 0.804 0.872 0.746 0.734
+[0.89;0.893] [0.289;0.294] [0.848;0.852] [0.888;0.892] [0.848;0.852] [0.801;0.808] [0.869;0.875] [0.742;0.751] [0.731;0.738]
+LGBM 0.917 0.212 0.881 0.94 0.881 0.851 0.916 0.794 0.797
+[0.915;0.918] [0.21;0.214] [0.879;0.883] [0.938;0.941] [0.879;0.883] [0.848;0.853] [0.913;0.918] [0.79;0.798] [0.793;0.801]
+HistGB 0.918 0.211 0.884 0.94 0.884 0.854 0.916 0.8 0.801
+[0.917;0.92] [0.209;0.213] [0.882;0.886] [0.939;0.942] [0.882;0.886] [0.851;0.857] [0.913;0.918] [0.796;0.803] [0.798;0.804]
+XGBoost 0.933 0.182 0.908 0.953 0.908 0.883 0.924 0.845 0.838
+[0.932;0.934] [0.18;0.184] [0.906;0.909] [0.952;0.954] [0.906;0.909] [0.881;0.885] [0.992;0.926] [0.842;0.848] [0.835;0.84]
+XGBoost 0.883 2.666 0.871 0.248 0.871 0.869 0.867 0.871 0.737
+RF 0.877 0.744 0.867 0.273 0.867 0.863 0.860 0.867 0.726
+GBM 0.857 1.476 0.854 0.42 0.854 0.843 0.836 0.854 0.689
+DT 0.843 21.54 0.827 0.332 0.827 0.824 0.822 0.827 0.649
+KNN 0.841 1.674 0.836 0.577 0.836 0.826 0.819 0.836 0.655
+AdaBoost 0.839 0.709 0.833 0.493 0.833 0.823 0.817 0.833 0.649
+SVC 0.837 0.641 0.849 0.332 0.849 0.826 0.818 0.849 0.666
+LR 0.802 12.24 0.817 0.362 0.817 0.791 0.785 0.817 0.601
 Table 4.
-Performance comparison of the $\text{Chrome}_{\textbf{22}}$-trained classifiers.
-Best results are bold, second best are underlined, and baseline models are in italic.
-The performance results of the t.ex-Graph classifier are reported in the bottom half (Raschke et al., 2023).
+Performance comparison of the $\text{Chrome}_{\textbf{22}}$-trained classifiers for identifying trackers across a subset of 13 metrics.
+The best results are highlighted in bold and blue, the second best are underlined and green, and baseline models are denoted as italic and orange.
+For comparison, the performance results of the T.EX-Graph classifier (Raschke et al., 2023) are reported in the bottom half of the table.
+6.1. Intra-browser Performance
 Figure 6.
-The radar charts illustrate that while all models perform well on the training set, the XGBoost model exhibits more consistent performance across the various test sets, suggesting better generalizability.
-The other models show reduced performance on test sets, indicating potential overfitting to the training data.
-In this section, we evaluate the effectiveness of our proposed methodology to detect trackers.
-We start by presenting our classifier results based on $\text{Chrome}_{22}$ data and follow-up by testing them with data from $\text{Firefox}_{22}$ and $\text{Brave}_{22}$.
-Furthermore, we assess the longitudinal performance with the $\text{Chrome}_{23}$ dataset that was collected almost a year later.
-We test our approach on an Apple computer, namely MacBook Pro (14-inch, 2021) with a 10-core Central Processing Unit (CPU), a 16-core Graphical Processing Unit (GPU), 1 Terabyte (TB) of SSD storage, and 32 Gigabyte (GB) of Random Access Memory (RAM).
-6.1. Intra-browser Performance
-Table 4 shows the performance of our $\text{Chrome}_{\textbf{22}}$-trained classifiers and the baselines.
-In terms of Accuracy, Ensemble Trees (ET) lead with a score of 0.964, closely followed by Random Forest (RF) and Decision Trees (DT) with scores of 0.960 and 0.956, respectively.
-The Log-Loss metric, which penalizes classifiers for incorrect predictions, shows that ET, RF, and DT also perform well, with ET having the lowest log loss, indicating a high level of confidence in its predictions.
-On the other hand, Gaussian Naive Bayes (GNB) exhibits a significantly higher log loss, which suggests that it is less confident in its predictions compared to the tree-based methods.
-The ROC-AUC metric further strengthens the position of ET, RF, and DT, as they all score above 0.94, indicating that they are good at differentiating between the positive and negative classes.
-The high MCC scores corroborate the findings from the other metrics, placing these three classifiers as the most robust performers across the board.
-In contrast, baseline models like LR are comparatively lower performance across these metrics.
-The gradient boosting models, particularly XGBoost, have fairly high scores as reflected in an AUC-ROC score of 0.915 and an F1-Score of 0.875.
-They are able to handle the class imbalance due to the fairly high MCC and AUPRC scores.
-However, their precision and recall scores indicate room for improvement in identifying instances of the minority class.
-Comparing the best performing classifiers ET and RF to the baseline models, shows that they perform significantly higher, except for the DT and its Log-Loss score.
+Radar charts plotting the performance of three selected classifiers across a subset of 13 metrics, illustrating that, while all models perform well on the in-distribution test set, the ET model exhibits more consistent performance across the various test sets, suggesting better generalizability.
+The models show a reduced performance on out-of-distribution test sets, indicating potential overfitting to the training data.
+Table 4 shows the performance of our $\text{Chrome}_{22}$-trained classifiers and the baselines.
+Regarding AUPRC, the RF leads with a score of 0.98, closely followed by Extra Trees (ET) classifier and DT with scores of 0.977 and 0.965, respectively.
+Except for the AdaBoost and GNB, all classifiers show a lower than baseline log-loss value, indicating a high prediction accuracy.
+The F1-score metric further strengthens the position of ET and RF, as they all score above 0.93, indicating that they are good at differentiating between the positive and negative classes (RQ2).
+The high MCC scores corroborate the other metrics, placing ET, RF, and DT as the most robust performers with accurate predictions across all categories in the confusion matrix.
+In contrast, baseline models like LR and AdaBoost perform comparatively lower across all metrics.
+The gradient boosting models, specifically XGBoost, have fairly high scores, as reflected in an AUPRC score of 0.953 and an F1-Score of 0.883.
+However, their MCC scores were lower in contrast to ET, RF, and DT due to their recall scores.
+Comparing the best-performing classifiers ET and RF to the baseline models shows that they perform significantly higher, except for DT, which only performed significantly worse in terms of log-loss.
 The t.ex-Graph classifiers display lower performance across all metrics.
 6.2. Cross-browser and Longitudinal Performance
-Figure 5 presents the cross-browser performance and longitudinal performance of our $\text{Chrome}_{\textbf{22}}$-trained classifiers.
-We focus on three selected classifiers, one baseline model, one ensemble tree-based model, and one ensemble tree-based model with gradient boosting.
-For Logistic Regression and Random Forest, $\text{Chrome}_{\textbf{22}}$ shows high performance across most metrics, which is expected as the models would be most adapted to the data they were trained on.
-The performance on $\text{Chrome}_{\textbf{23}}$, $\text{Firefox}_{\textbf{22}}$, and $\text{Brave}_{\textbf{22}}$ varies, with a general trend of decreasing performance, indicating potential overfitting to the training data or a lack of generalizability to other datasets.
-The XGBoost model shows a similar trend, with high performance on the training set and slightly varied results on the test sets.
-However, the variations in the XGBoost model’s performance are less pronounced than in the Logistic Regression and Random Forest models, suggesting that XGBoost may be more robust or generalizable across different browser datasets.
-Overall, the figure indicates that while all models perform well on the dataset they were trained on, their ability to generalize to other data is varied.
-XGBoost appears to have the most consistent performance across different browser datasets, indicating a potential advantage in terms of generalization.
+Figure 6 presents the cross-browser performance and longitudinal performance of our $\text{Chrome}_{22}$-trained classifiers.
+We focus on three selected classifiers: one baseline model, our best performing model, and one gradient boosting model.
+For ET and XGBoost, $\text{Chrome}_{22}$ shows high performance across most metrics, which is expected as the models would be most adapted to the data they were trained on.
+The performance on $\text{Chrome}_{23}$, $\text{Firefox}_{22}$, and $\text{Brave}_{22}$ varies, with a general trend of decreasing performance, indicating potential overfitting to the training data or a lack of generalizability to other datasets.
+The former should not be the case, as we observe less than a 5% performance drop between the in-distribution training and test set.
+Furthermore, we split the in-distribution data before the pre-processing, thus mitigating any data leakage and applied Repeated Stratified K-Fold cross-validation with five repeats and five folds.
+As shown in the previous sections, the datasets are different and trackers have varying structures, which is why we perform a cross-browser evaluation to test the generalizability.
+Overall, the results depicted in Figure 6 indicate that, while all models perform well on the dataset they were trained on, their ability to generalize to other data varies.
+ET exhibits the most consistent performance across different browser datasets, indicating better generalization but it had a lower log-loss score for $\text{Firefox}_{22}$ compared to the two.
 The figure underscores the importance of cross-browser testing to ensure that models are not only tuned to a specific dataset but also maintain their performance across diverse scenarios.
-The Brave results show a high accuracy metric for the overall performance, but the F1-Score and MCC are low and represent a more realistic performance of the classifiers, which is expected due to the more considerable differences between the Chrome and Brave data.
-The longitudinal performance analysis, as depicted in Figure 5, indicates a slight decline in the efficacy of the classification models when transitioning from the $\text{Chrome}_{\textbf{22}}$ to the $\text{Chrome}_{\textbf{23}}$ dataset.
-This is evidenced by the contraction of the metric lines from $\text{Chrome}_{\textbf{22}}$ to $\text{Chrome}_{\textbf{23}}$ across the three models, suggesting a decrease in metrics such as AUC-ROC, Accuracy, and others.
-Notably, the XGBoost model demonstrates the smallest degree of performance reduction, underscoring its robustness and adaptability to changes over time.
-In contrast, both the Logistic Regression and Random Forest models exhibit a more pronounced drop, highlighting potential challenges in model generalization across browser version updates.
-This suggests that while all models are affected by evolving data characteristics, ensemble methods, particularly gradient boosting as implemented in XGBoost, may offer a more stable predictive performance in longitudinal settings.
-Applying isotonic calibration had only minor improvements compared to the uncalibrated models as shown in Figure 6.
+The Brave results show a high accuracy metric for the overall performance, but the F1-Score and MCC are low and represent a more realistic performance, which is expected due to the considerable differences between the Chrome and Brave data.
+Our longitudinal performance analysis indicates a slight decline in the efficacy of the classifiers when transitioning from the $\text{Chrome}_{22}$ to the $\text{Chrome}_{23}$ dataset.
+We can observe a decrease in metrics such as AUC-ROC, Accuracy, and others.
+Notably, the MCC score between ET and XGBoost are getting similar.
+A trend which we can observe across most classifiers for $\text{Chrome}_{23}$ but also $\text{Firefox}_{22}$.
+However, the MCC scores of the gradient boosting models did not drop significantly more between $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$, e.g., XGBoost went down from 0.838 to 0.823, whereas the ET suffered from 0.903 to 0.851.
+This suggests that while all models are affected by evolving data characteristics, ensemble methods, particularly some gradient boosting models, may offer a more stable predictive performance in longitudinal settings in contrast to the ET and RF.
+Nevertheless, ET and RF had better or similar performance compared to XGBoost.
+Applying isotonic calibration has only minor improvements compared to the uncalibrated models, as highlighted by Figure 9.
+6.3. Feature Importance
+As shown in Figure 7, the features Cross-Origin-Resource-Policy and Last-Modified were the important features overall, followed by Accept-Ranges and P3P.
+These four headers have distinct purposes in HTTP communication, except for P3P which is deprecated nowadays  [^8] https://www.w3.org/TR/P3P/.
+Although we plot the top ten features, the majority of them have low importance values below 0.1.
+DT, RF, and GBM share similar values, whereas the best-performing model, ET, has no features with a significantly higher value than the others.
+Figure 7.
+The top ten features are presented in order of importance across different classifiers based on the in-distribution test set.
+The heatmap illustrates the normalized importance values for each feature and classifier.
+We calculated Permutation feature importance for LR and GNB but not for the others as we do not have any high cardinality features.
+Iqbal et al.
+(Iqbal et al., 2022), who also used response header data for their classification as discussed in Section 2.1, calculated the information gain for each of their eight response-based features.
+This is only partially comparable to feature importance but enables two key comparisons: (i) ETag and especially P3P were important features as well, and (ii) they highlighted the importance of the Content-Length, Content-Type, and Set-Cookie headers, although they did not rank among the top 10 features in our classifiers.
+6.4. HTTP Request Header Performance
+Lastly, we compare the performance of our response header-based approach to classifiers that were trained on request headers.
+Comparing the request-based classifiers internally, we observe performance differences similar to those of their response-based counterparts.
+Overall, however, the request-based classifiers perform worse, as shown in Figure 8.
+Their AUPRC, F1-Score, and MCC values are close to 0.5 or less, which means that their prediction reliability is low.
+Figure 8. Comparison between our response-based, request-based, and T.EX-Graph’s Random Forest classifier. The response-based classifier outperforms its counterparts across all evaluated metrics.
+We also observe a pattern similar to the response-based classifiers regarding cross-browser and longitudinal performance.
+Over time, the classifiers perform slightly worse, e.g., with the MCC for the ET going down from 0.42 to 0.404.
+Regarding $\text{Firefox}_{22}$ data, the precision improved but the recall decreased.
+Nevertheless, the MCC is only 0.213, and the AUPRC is 0.458 for the ET, indicating that the classifier has no discriminative power.
+In the case of Brave, the classifiers are essentially ineffective and exhibit a low F1-score and MCC.
+Among the most important features across the majority of classifiers are X-Client-Data, Content-Type, Origin, and Cookie.
+The first two are especially relevant, with values ranging from 0.2 to 0.566, excluding GNB, AdaBoost, and LightGBM.
+Although Content-Type is both a request and response header, it was not among the top-ten most important features for the response classifiers.
+The most important feature for the request-based classifiers, X-Client-Data, is used by Chrome to send an identifier to certain Alphabet-affiliated domains, e.g., doubleclick.net or googleadservices.com  [^9] https://github.com/chromium/chromium/blob/master/components/variations/net/variations_http_headers.cc, which are prevalent in our datasets.
+We observe similar importance values when we use an out-of-distribution dataset like $\text{Firefox}_{22}$.
 7. Discussion
-Tracking affects users across all three browsers. However, this varies to a certain degree, e.g. Brave had the least amount of tracking-related HTTP responses compared to Chrome and Firefox.
-Nevertheless, an imbalance can be observed in all data sets, which might be a problem for specific ML models.
-Even though many different methods were developed to balance data sets, due to the moderate imbalance in Chrome and Firefox and the good performance, a resampling was not conducted to preserve the observed distribution.
-Still, the large number of observations provides sufficient HTTP response messages for each class label.
-Furthermore, several evaluation metrics were calculated and provided that are aimed at imbalanced data sets, e.g.
-BACC.
-The classifiers performed similarly in Chrome and Firefox but were poor with Brave data.
-First and foremost, the application of binary-encoded features resulted in surprisingly good results, i.e.
-the mere presence of a header value is enough and extensive feature engineering is not required from a pure tracking detection standpoint.
-On the other hand, explaining the models is limited, except for model parameters, training time, or the calculated metrics.
-In addition, this method simplifies the models to a high degree, although HTTP header values can reveal more information about trackers and their differences from non-trackers.
-Nevertheless, the binary approach helped preserve the core information, namely whether a header is related to tracking.
-The NB classifier did not perform as well as others, but that might be due to its limited applicability to the data sets.
-As mentioned above, NB is based on Bayes Theorem and assumes that features within a class are not correlated to each other.
-It can be assumed that a set of headers and thus features are correlated to a certain degree.
-These were not removed which is a violation of Bayes Theorem in a statistical sense could be argued, and thus, its applicability.
-On the contrary, a paper by Berrar (Berrar, 2018) argues that NB can perform well, even under these circumstances.
-Other implementations, e.g., the Categorical Naïve Bayes might have been a good fit for the training sets with categorical features (nb, 2023).
-Filter lists are reliable and effective when it comes to blocking trackers (Mazel et al., 2019), and are one of the oldest methods, e.g.
-EasyList was created in 2005 (eas, 2023a).
-They present a popular ground truth to label data sets for training ML models.
-However, they have to be constantly updated, which commonly happens manually.
+Tracking affects users across all three browsers but varies to a certain degree, e.g., Brave had the least amount of tracking-related HTTP messages compared to Chrome and Firefox.
+This can be attributed to Brave’s extensive built-in privacy protection mechanisms.
+First and foremost, the application of binarization resulted in surprisingly good results, i.e., the mere presence of a header value is sufficient and extensive feature engineering is not required from a tracker detection standpoint.
+Our approach works well in this imbalanced setting, although it might be problematic for some classification algorithms without additional measures, such as setting class weights or rebalancing the class distribution.
+Additional measures include using ML models, which handle imbalanced data sets well (Juez-Gil et al., 2021), or anomaly detection (Park et al., 2018).
+Thus, although this has not been considered in our pipeline, it might be interesting to explore how well these measures may improve the performance of our proposed approach.
+Spoofing of header values to circumvent our approach should not be a problem, as the values are not taken into account.
+However, this limits the models’ explainability in the sense that we lose the actual header values that might provide us with relevant information to understand trackers.
+Servers can modify a response by adding or removing headers before responding to a client.
+The latter could impact the effectiveness of our classifiers, as they require the presence of headers.
+Because analyzing the characteristics and use cases of each header in our datasets is unfeasible, we focus on the top ten headers from our feature importance analysis to assess the impact of their removal on our classifiers’ ability to correctly identify trackers.
+For this, we must differentiate between deprecated and actively used headers.
+The former may be removed by servers at any time and might prevent a positive detection by our classifiers.
+Examples include the Pragma header, which provides caching directives and has been superseded by the Cache-Control header, and the P3P header representing privacy practices.
+Headers belonging to the latter category cannot be removed as easily.
+Security headers, such as X-Frame-Options, Cross-Origin-Resource-Policy, and Content-Security-Policy-Report-Only, are used to defend against various attacks like click-jacking, cross-site script inclusion attacks, or cross-site request forgery.
+Of equal importance are headers related to Cross-Origin-Resource-Sharing, such as Access-Control-Allow-Credentials that allows the inclusion of, e.g., credentials for HTTP authentication or cookies in cross-origin requests.
+Other vital headers include Last-Modified, Accept-Ranges, Etag, and Content-Encoding, which are relevant for compression, caching, or downloads.
+Excluding any of the aforementioned headers would have consequences leading to decreased security, performance, or functionality.
+Therefore, it would not be in the interest of web services to exclude these headers.
+Our cross-browser approach reveals three things: firstly, the $\text{Chrome}_{22}$ dataset already captures most of the tracker structure.
+Secondly, due to the resulting effectiveness of a $\text{Chrome}_{22}$-trained classifier on out-of-distribution data, such as $\text{Firefox}_{22}$ and $\text{Chrome}_{23}$, it may not be necessary to re-train classifiers for these datasets.
+Thirdly, applying our approach to browsers like Brave that filter web tracking by default has a strong negative performance impact because the classifiers trained on unfiltered data are not able to effectively classify the modified structure of trackers in the filtered data.
+The results of our clustering shown in Figure 5 reveal that trackers are more dispersed than non-trackers, which might be an argument for browser-specific models, as these might be able to better distinguish between trackers.
+Another challenge is the overlap between trackers and non-trackers, which further impedes correct classification.
+Feature engineering taking the header values into account might be a valuable approach to reduce the overlap, as header values can differ between trackers and non-trackers (see Section 4).
+A data-driven and large-scale in-depth analysis of header values and tracker structures across browsers is needed to explore this approach further.
+Our comparison of request- and response-based classifiers revealed that request-based classifiers perform worse overall, which might be attributed to their comparatively small feature space, i.e., a smaller number of unique headers.
+This weaker performance suggests that request headers are less informative for the detection of web trackers than response headers.
+Nevertheless, the request-based classifiers correctly identify some trackers that the response-based classifiers missed, and vice-versa, as highlighted by the example of ET in Table 5.
+Thus, future work could explore combinations of both approaches as part of a multi-modal ensemble model.
+Predicted Label by ET True Label Count Total
+Response Request
+NT NT T 3777 103357
+NT NT NT 99580 103357
+T T NT 404 7826
+T T T 7422 7826
+NT T T 729 19257
+NT T NT 18528 19257
+T NT NT 1991 41304
+T NT T 39313 41304
+Table 5. Comparison of the predicted and true labels for the request- and response-based ET classifier based on the in-distribution test set.
+Utilizing the models for automated web tracker detection in real-life application scenarios is possible in two scenarios.
+First, HTTP response information is unavailable in an online (or live) detection scenario to block communication with web trackers.
+In this scenario, our model can enable or support reinforced learning of a complementary model on whose prediction blocking of HTTP communication is decided.
+This process, e.g., may imply manual inspection of HTTP communication where the predictions of our model are considered in cases where the primary model is wrong or indecisive.
+Second, as filter lists are still fundamental for web privacy protection, the manual process of generating filter rules can be optimized or outright replaced by utilizing our model.
+In this scenario, existing data sets or data from volunteers containing HTTP communication are used to identify web tracking requests.
+Filter rules for those can be dynamically generated and appended to conventional filter lists, which are already used.
+As one of the oldest methods for blocking trackers (Mazel et al., 2019), filter lists have proven to be reliable and effective.
+However, they have to be updated constantly, which is commonly done manually.
 A longitudinal study of nine years by Alrizah et al.
-(Alrizah et al., 2019) showed that human errors, e.g.
-acting in a timely manner or incorrect reports of possible new trackers and ads, are a root cause for lower performance.
+(Alrizah et al., 2019) showed that human errors and delayed updates are a root cause for lower performance.
 Chen et al.
-(Chen et al., 2021) highlighted evasion methods to circumvent blockers and filter lists: (i) changing URLs, (ii) embedding code directly into the website with the script tag, and (iii) combining functional and tracking code in one file.
-In summary, using filter lists for our ground truth is not without problems and directly affect the performance of our classifiers.
-7.1. Limitations
-Firstly, the analysis of response headers to fill the identified research gap has its strengths as well as its weaknesses.
-Headers represent only a part of an HTTP message, and research has shown that HTTP bodies are also relevant when it comes to understanding and identifying trackers (Li et al., 2015; Guarino.
+(Chen et al., 2021) highlighted evasion methods to circumvent blockers and filter lists: (i) changing URLs, (ii) embedding code directly into the website with the script-tag, and (iii) combining functional and tracking code in one file.
+In summary, using filter lists for our ground truth is not without problems and may impact the evaluation of our classifiers.
+Nevertheless, they present a popular ground truth in research to label datasets for training ML models.
+8. Conclusion
+In this work, we train $\text{Chrome}_{22}$-based classifiers on a subset of response headers to evaluate their effectiveness and usefulness in web tracker detection while retaining the natural imbalance of trackers.
+Our two best classifiers RF and ET achieve scores ranging from 0.9 to 0.98.
+In addition, we measure the cross-browser effectiveness, showing that multiple $\text{Chrome}_{22}$-based classifiers achieve high AUPRC but moderate MCC values for $\text{Firefox}_{22}$.
+However, we also observe large performance losses for $\text{Brave}_{22}$.
+The longitudinal performance of the selected ML models deteriorates only by small margins, which highlights their potential due to the changing capabilities of trackers (Fouad et al., 2018; Bujlow et al., 2017; Englehardt and Narayanan, 2016).
+Still, periodic re-training might be necessary if the performance decreases significantly.
+We observe that response headers represent characteristics of web trackers which were previously under-explored by past research.
+They can be used to differentiate well between trackers and non-trackers and provide additional information on the nature of tracking-related entities.
+Our data-driven approach combined with binarization reveales new and confirms previously identified features.
+Applying the same methodology to request headers results in a weaker performance, thus strengthening the feasibility of using response headers for web tracker detection.
+Moreover, headers represent only a part of an HTTP message, and research has shown that the message bodies are also relevant when it comes to understanding and identifying trackers (Li et al., 2015; Guarino.
 et al., 2020).
-From a broader point of view, only one part of application layer traffic is analyzed, disregarding other parts of web traffic, e.g.
-transport layer data.
-Secondly, the data set could have been processed differently across the pipeline.
-As above-mentioned, the data sets are imbalanced, with a 72:28 non-tracker-to-tracker ratio for Chrome and Firefox data.
-It is not a severe form of imbalance, but it might be problematic for some classification algorithms.
-They could be biased against the majority class, and it might be possible to consistently predict the non-tracker label and still observe good metrics.
-Several resampling techniques exist, simple methods, e.g.
-over- and undersampling (Hernandez et al., 2013), or more complex methods, e.g.
-SMOTE (Chawla et al., 2002), or one of its variations (Han et al., 2005; Mukherjee and Khushi, 2021).
-Additional measures include using ML models, which handle imbalanced data sets well (Juez-Gil et al., 2021), or anomaly detection (Park et al., 2018).
-In summary, this aspect of the data pipeline has not been implemented with any of the aforementioned methods, and it might be interesting to explore how well these techniques may work for the present concept.
-Thirdly, the labeling of the initial data sets, i.e.
-the quality of the ground truth.
-T.EX allows the implementation of different filter lists and offers three to date, namely EasyList, EasyPrivacy, and Disconnect.me.
-All base their functionality on extensive filter lists, i.e.
-rules and RegEx are used to block tracking-related HTTP requests.
-Updates are carried out regularly every four days for the first two (eas, 2023a, b), and on an irregular schedule for the latter (dis, 2023).
-Research has shown that these filter lists are not without errors regarding their rules (Snyder et al., 2020).
-8. Conclusion
-In this work, we trained Chrome-based classifiers on a subset of binary-encoded HTTP/S response headers to evaluate their effectiveness and usefulness in web tracker detection while retaining the natural imbalance of trackers.
-In addition, we measured the cross-browser effectiveness, showing that multiple Chrome-based classifiers achieve high accuracies for Firefox as well.
-However, large performance losses were observed for Brave.
-The longitudinal performance of the selected ML models deteriorated only by small margins, which highlights their potential due to the changing capabilities of trackers (Fouad et al., 2018; Bujlow et al., 2017; Englehardt and Narayanan, 2016).
-Response headers can be used to differentiate well between trackers and non-trackers and they provide additional information to the nature of tracking-related entities.
-For our future research we plan to further investigate response headers, focusing on the values for feature engineering, or the use of Deep Learning models.
-Lastly, we will approach the deployment aspect two-fold, one from a user perspective in form of browser extension to test the models in a real-life scenario, and secondly as a fully automated system in the cloud that integrates a crawler to periodically update the models.
+Emerging technologies like Large Language Models could be used to analyze bodies in an efficient and meaningful way.
+In future research, we plan to expand on our approach of leveraging response headers for web tracker detection by considering header values in feature engineering and exploring the viability of deep-learning approaches based on response header data.
+Additionally, we intend to integrate our approach with existing solutions to develop a holistic web tracking detection system.
+Acknowledgements.
+We thank our anonymous reviewers for their feedback.

### v3

@@
 (Wu et al., 2016) developed DMTTrackerDetector, a system using decision trees in which observed API calls serve as features.
-Their system, comprising a 717-dimensional feature vector, achieved 97.8% accuracy and addressed obfuscation issues, showing the distinct JavaScript API usage by trackers and non-trackers.
+Their system, comprising a 717-dimen- sional feature vector, achieved 97.8% accuracy and addressed obfuscation issues, showing the distinct JavaScript API usage by trackers and non-trackers.
 Bahrami et al.
@@
 Figure 1. Frequency distribution of headers for each browser, indicating the frequency with which each header is present in the dataset.
+[…]
+[Frequency distribution of headers for each browser, indicating the frequency with which each header is present in the dataset.]In this frequency plot we can observe that only a few number of headers are prevalent and most headers are rarely seen in the wild.
 Figure 1 shows a frequency plot to outline how often each header appears in the dataset.
@@
 Venn diagram of HTTP response headers in the examined datasets, revealing a core set of 5672 headers that are common across Chrome, Firefox, and Brave.
+Unique headers are present in each browser, although $\text{Chrome}_{23}$ has the highest number of unique headers, which may reflect browser-specific features or experimental headers.
+[…]
+[Venn diagram of HTTP response headers.]Venn diagram of HTTP response headers in the examined datasets, revealing a core set of 5672 headers that are common across Chrome, Firefox, and Brave.
 Unique headers are present in each browser, although $\text{Chrome}_{23}$ has the highest number of unique headers, which may reflect browser-specific features or experimental headers.
@@
 Row (A) presents the ECDFs for both Chrome datasets for the Content-Length header values.
+We set a cut-off value at 10,000 to highlight the core observation and more than half of the responses have a value below this threshold – $\tilde{\mu}_{T22}=62$, $\tilde{\mu}_{NT22}=8068$ and $\tilde{\mu}_{T23}=45$, $\tilde{\mu}_{NT23}=8021.5$.
+Row (B) shows how similar the values across $\text{Chrome}_{22}$, $\text{Firefox}_{22}$, and $\text{Brave}_{22}$ are for the X-XSS-Protection header.
+[…]
+[ECDF plots for two selected response headers.]Row (A) presents the ECDFs for both Chrome datasets for the Content-Length header values.
 We set a cut-off value at 10,000 to highlight the core observation and more than half of the responses have a value below this threshold – $\tilde{\mu}_{T22}=62$, $\tilde{\mu}_{NT22}=8068$ and $\tilde{\mu}_{T23}=45$, $\tilde{\mu}_{NT23}=8021.5$.
@@
 Therefore, applying a process that achieves a reduced feature space $X_{i}^{\prime}\subseteq X$ is necessary, entailing a reasonable training time.
-  • (CH5) Data Similarity: The datasets collected across browsers can exhibit pronounced similarities in their feature distributions and values.
+  • (CH5) Data Similarity: The datasets collected across brow- sers can exhibit pronounced similarities in their feature distributions and values.
 Given the high dimensionality of our feature space, exhaustive pairwise comparisons can be computationally expensive and time-consuming, highlighting the challenge of accurately gauging data similarities in such a complex space.
 Figure 4. The parameterized pipeline automatically processes the collected datasets and trains the classifiers.
+[…]
+[The parameterized pipeline automatically processes the collected datasets and trains the classifiers.]The parameterized pipeline automatically processes the collected datasets and trains the classifiers.
+The pipeline starts with data collection and labeling.
+After the data is initially parsed, it is prepared and processed.
+Lastly, the models are trained and tested, and calibrated.
 5.2. Machine Learning Pipeline
@@
 We also compute metrics that account for or are unaffected by imbalance, and the selected ensemble models are inherently robust to this issue.
-Lastly, following the findings of Niculescu-Mizil and Caruana (Niculescu-Mizil and Caruana, 2005), we apply Isotonic Regression instead of Platt Scaling to calibrate the probabilities calculated by our models.
+Last, following the findings of Niculescu-Mizil and Caruana (Niculescu-Mizil and Caruana, 2005), we apply Isotonic Regression instead of Platt Scaling to calibrate the probabilities calculated by our models.
 This is done by fitting an Isotonic Regression model to the models’ predictions as well as the true labels to adjust the predicted probabilities.
@@
 However, accuracy is not useful with imbalanced datasets because a model could always predict the majority class and achieve accuracy with the same value as the distribution of said class.
-For this reason, we calculate the following additional metrics to quantify the performance of a classifier with an imbalanced dataset:
-Balanced Accuracy:
-Broderson et al.
-(Brodersen et al., 2010) introduced the Balanced Accuracy (BACC) metric for imbalanced datasets in 2010 as an alternative to average accuracy scores, which are, according to the authors, problematic for CIs and imbalanced datasets.
-The metric comprises recall and true negative rate (TNR) and is defined as:
-(2) $BalancedAccuracy=\frac{1}{2}\left(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\right)$
-Log-loss Score:
-Also referred to as cross-entropy-loss, the log-loss score focuses on the predictive performance of a model.
-It describes how large the error between a prediction probability and the true label is, where a value closer to zero represents a better score.
-However, this metric can be influenced by an imbalance, i.e., the baseline log loss gets smaller with an increase in imbalance.
-Therefore, it is essential to compare the metric to its baseline value.
-For instance, the baseline for our $\text{Chrome}_{22}$ dataset with a 70:30 ratio equals 0.611  [^7] https://towardsdatascience.com/intuition-behind-log-loss-score-4e0c9979680a.
-The model’s log-loss should be lower than its baseline.
-The log-loss score is defined as:
-(3) $L_{log}(y,p)=-[y\cdot log(p)+(1-y)\cdot log(1-p)]$
-where $y$ is the label class and $p$ is the probability.
-Matthews-Correlation-Coefficient:
-The Matthews-Correlation-Coefficient (MCC) (Matthews, 1975), a metric that has seen increasing popularity in recent years, is defined as:
-(4) $MCC=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)\cdot(TP+FN)\cdot(TN+FP)\cdot(TN+FN)}}$
-MCC ranges from -1 to +1 and can be understood as a correlation coefficient between the true and predicted labels.
-The MCC will only achieve a high value if both class labels are predicted with high accuracy, positioning the MCC as a meaningful metric for imbalanced datasets (Boughorbel et al., 2017).
-This was further substantiated in 2020 by Chicco et al.
-(Chicco and Jurman, 2020), who outlined how accuracy and F1-Score are not reliable measures for imbalanced datasets.
-AUPRC:
-The last metric in our study is the Area Under the Precision-Recall Curve (AUPRC).
-This metric is relevant when dealing with imbalanced datasets.
-The Precision-Recall Curve is a plot that illustrates the trade-off between precision and recall for a classifier at various thresholds.
-The AUPRC is an aggregated measure of this curve, representing the classifier’s ability to correctly identify positive instances across different thresholds.
-In imbalanced settings, where one class is significantly underrepresented, AUPRC is a more informative and reliable metric than ROC-AUC, as it focuses on the performance of the classifier on the minority class, which is of interest for us (Saito and Rehmsmeier, 2015).
+For this reason, we calculate the following additional metrics to quantify the performance of a classifier with an imbalanced dataset: (i) Balanced Accuracy (BACC), (ii) Log-loss Score, (iii) Matthews-Correlation-Coefficient (MCC), and (iv) Area Under the Precision-Recall Curve (AUPRC).
+A further explanation for each metric is reported in the Appendix B.
 6. Experiments
@@
 The results are separated by class to highlight differences between trackers and non-trackers.
-In this section, we evaluate the effectiveness of our proposed methodology in detecting trackers.
+[…]
+[t-SNE plots across browsers to gauge data similarity.]t-SNE plots using a representative sample across all browsers to gauge data similarity.
+Separate clusters exist for trackers and non-trackers, but there is significant overlap between both classes.
+The results are separated by class to highlight differences between trackers and non-trackers.
+In this section, we evaluate the effectiveness of our proposed metho- dology in detecting trackers.
 We begin by presenting our classifier results based on $\text{Chrome}_{22}$ data and follow up by testing them with data from $\text{Firefox}_{22}$ and $\text{Brave}_{22}$.
 Furthermore, we assess the longitudinal performance with the $\text{Chrome}_{23}$ dataset that was collected almost a year after $\text{Chrome}_{22}$.
-Note that this section only summarizes our results; performance metrics across all classifiers are reported in the Appendix 9.1 and 9.2.
+Note that this section only summarizes our results.
+Performance metrics across all classifiers are reported in the Appendix D.
 We test our approach on a MacBook Pro (14-inch, 2021) with a 10-core CPU, a 16-core GPU, 1TB of SSD storage, and 32GB of RAM.
@@
 The models show a reduced performance on out-of-distribution test sets, indicating potential overfitting to the training data.
+[…]
+[Radar charts of the LR, ET, and XGBoost classifier for cross-browser performance.]Radar charts plotting the performance of three selected classifiers across a subset of 13 metrics, illustrating that, while all models perform well on the in-distribution test set, the ET model exhibits more consistent performance across the various test sets, suggesting better generalizability.
+The models show a reduced performance on out-of-distribution test sets, indicating potential overfitting to the training data.
 Table 4 shows the performance of our $\text{Chrome}_{22}$-trained classifiers and the baselines.
@@
 Applying isotonic calibration has only minor improvements compared to the uncalibrated models, as highlighted by Figure 9.
+The exceptions are GNB and AdaBoost, where estimates have improved, e.g., the uncalibrated AdaBoost initially centered most predictions around 0.5.
 6.3. Feature Importance
@@
 The top ten features are presented in order of importance across different classifiers based on the in-distribution test set.
+The heatmap illustrates the normalized importance values for each feature and classifier.
+We calculated Permutation feature importance for LR and GNB but not for the others as we do not have any high cardinality features.
+[…]
+[Feature importance of Chrome-based classifiers using HTTP response headers.]The top ten features are presented in order of importance across different classifiers based on the in-distribution test set.
 The heatmap illustrates the normalized importance values for each feature and classifier.
@@
 Their AUPRC, F1-Score, and MCC values are close to 0.5 or less, which means that their prediction reliability is low.
-Figure 8. Comparison between our response-based, request-based, and T.EX-Graph’s Random Forest classifier. The response-based classifier outperforms its counterparts across all evaluated metrics.
+Figure 8. Comparison between our response-based, request-based, and T.EX-Graph’s RF classifier. The response-based classifier outperforms its counterparts across all evaluated metrics.
+[…]
+[The RF classifier outperforms the baselines.]Comparison between our response-based, request-based, and T.EX-Graph’s RF classifier.
+The response-based classifier outperforms its counterparts across all evaluated metrics.
 We also observe a pattern similar to the response-based classifiers regarding cross-browser and longitudinal performance.
@@
 This can be attributed to Brave’s extensive built-in privacy protection mechanisms.
+7.1. Experimental Results
 First and foremost, the application of binarization resulted in surprisingly good results, i.e., the mere presence of a header value is sufficient and extensive feature engineering is not required from a tracker detection standpoint.
@@
 Thus, although this has not been considered in our pipeline, it might be interesting to explore how well these measures may improve the performance of our proposed approach.
-Spoofing of header values to circumvent our approach should not be a problem, as the values are not taken into account.
-However, this limits the models’ explainability in the sense that we lose the actual header values that might provide us with relevant information to understand trackers.
-Servers can modify a response by adding or removing headers before responding to a client.
-The latter could impact the effectiveness of our classifiers, as they require the presence of headers.
-Because analyzing the characteristics and use cases of each header in our datasets is unfeasible, we focus on the top ten headers from our feature importance analysis to assess the impact of their removal on our classifiers’ ability to correctly identify trackers.
-For this, we must differentiate between deprecated and actively used headers.
-The former may be removed by servers at any time and might prevent a positive detection by our classifiers.
-Examples include the Pragma header, which provides caching directives and has been superseded by the Cache-Control header, and the P3P header representing privacy practices.
-Headers belonging to the latter category cannot be removed as easily.
-Security headers, such as X-Frame-Options, Cross-Origin-Resource-Policy, and Content-Security-Policy-Report-Only, are used to defend against various attacks like click-jacking, cross-site script inclusion attacks, or cross-site request forgery.
-Of equal importance are headers related to Cross-Origin-Resource-Sharing, such as Access-Control-Allow-Credentials that allows the inclusion of, e.g., credentials for HTTP authentication or cookies in cross-origin requests.
-Other vital headers include Last-Modified, Accept-Ranges, Etag, and Content-Encoding, which are relevant for compression, caching, or downloads.
-Excluding any of the aforementioned headers would have consequences leading to decreased security, performance, or functionality.
-Therefore, it would not be in the interest of web services to exclude these headers.
 Our cross-browser approach reveals three things: firstly, the $\text{Chrome}_{22}$ dataset already captures most of the tracker structure.
-Secondly, due to the resulting effectiveness of a $\text{Chrome}_{22}$-trained classifier on out-of-distribution data, such as $\text{Firefox}_{22}$ and $\text{Chrome}_{23}$, it may not be necessary to re-train classifiers for these datasets.
+Secondly, due to the resulting effectiveness of a $\text{Chrome}_{22}$-trained classifier on out-of-distribution data, such as $\text{Firefox}_{22}$, it may not be necessary to re-train classifiers for these datasets.
 Thirdly, applying our approach to browsers like Brave that filter web tracking by default has a strong negative performance impact because the classifiers trained on unfiltered data are not able to effectively classify the modified structure of trackers in the filtered data.
@@
 Table 5. Comparison of the predicted and true labels for the request- and response-based ET classifier based on the in-distribution test set.
-Utilizing the models for automated web tracker detection in real-life application scenarios is possible in two scenarios.
+7.2. Potential Evasion Approaches
+Spoofing of header values to evade our approach should not be a problem, as the values are not taken into account.
+However, this limits the models’ explainability in the sense that we lose the actual header values that might provide us with relevant information to understand trackers and whether some headers are required by trackers or not.
+Servers can modify a response by adding or removing headers before responding to a client.
+The latter could impact the effectiveness of our classifiers, as they require the presence of headers.
+Because analyzing the characteristics and use cases of each header in our datasets is unfeasible, we focus on the top ten headers from our feature importance analysis to assess the impact of their removal on our classifiers’ ability to correctly identify trackers.
+For this, we must differentiate between deprecated and actively used headers.
+The former may be removed by servers at any time and might prevent a positive detection by our classifiers.
+Examples include the Pragma header, which provides caching directives and has been superseded by the Cache-Control header, and the P3P header representing privacy practices.
+Headers belonging to the latter category cannot be removed as easily.
+Security headers, such as X-Frame-Options, Cross-Origin-Resource-Policy, and Content-Security-Policy-Report-Only, are used to defend against various attacks like click-jacking, cross-site script inclusion attacks, or cross-site request forgery.
+Of equal importance are headers related to Cross-Origin-Resource-Sharing, such as Access-Control-Allow-Credentials that allows the inclusion of, e.g., credentials for HTTP authentication or cookies in cross-origin requests.
+Other vital headers include Last-Modified, Accept-Ranges, Etag, and Content-Encoding, which are relevant for compression, caching, or downloads.
+Excluding any of the aforementioned headers would have consequences leading to decreased security, performance, or functionality.
+Therefore, it would not be in the interest of web services to exclude these headers.
+Siby et al.
+(Siby et al., 2022) argued that content features, such as URLs, are more vulnerable to obfuscation than structural features.
+This assertion extends to the content of headers as well.
+However, our classifiers are not concerned with the content of headers, but rather with their mere presence.
+Malicious actors could employ two different techniques in this case: (i) the addition of headers and (ii) the removal of headers – both could be done intentionally or randomly.
+The complete removal, however, could potentially disrupt the functionality of trackers or non-trackers.
+The presented classifiers could be further enhanced through the integration or the extension of structural and flow features.
+In the case of Siby et al., the inclusion of these features contributed to the robustness of AdGraph.
+The authors (Siby et al., 2022) developed a threat model and considered various techniques that trackers might utilize to evade detection.
+Although it is conceivable that an individual tracker could obfuscate its identity by employing a particular set of response headers to mimic non-trackers, if all trackers were to adopt a uniform set of headers or employ a similar strategy, they would collectively form a distinct cluster as illustrated in Figure 5, thereby differentiating themselves from non-trackers.
+Furthermore, high coordination between trackers would be necessary to employ said strategies but due to the heterogeneity of trackers and their owners, this level of coordination is unlikely.
+Nevertheless, a large impact could be achieved if the owner’s trackers are highly prevalent such as Alphabet’s.
+Lastly, the similarity between trackers and non-trackers with respect to their feature vectors is already notably, emphasizing the classifier’s effectiveness in differentiating between these two groups.
+This suggests that the classifier is already highly proficient in identifying and separating trackers from non-trackers.
+7.3. Deployment Scenarios
+Utilizing the models for automated web tracker detection in real-life applications is possible in two scenarios.
 First, HTTP response information is unavailable in an online (or live) detection scenario to block communication with web trackers.
@@
 Filter rules for those can be dynamically generated and appended to conventional filter lists, which are already used.
+7.4. Limitations of Using Filter Lists
 As one of the oldest methods for blocking trackers (Mazel et al., 2019), filter lists have proven to be reliable and effective.
@@
 Chen et al.
-(Chen et al., 2021) highlighted evasion methods to circumvent blockers and filter lists: (i) changing URLs, (ii) embedding code directly into the website with the script-tag, and (iii) combining functional and tracking code in one file.
+(Chen et al., 2021) highlighted methods to evade blockers and filter lists: (i) changing URLs, (ii) embedding code directly into the website with the script-tag, and (iii) combining functional and tracking code in one file.
 In summary, using filter lists for our ground truth is not without problems and may impact the evaluation of our classifiers.
@@
 8. Conclusion
-In this work, we train $\text{Chrome}_{22}$-based classifiers on a subset of response headers to evaluate their effectiveness and usefulness in web tracker detection while retaining the natural imbalance of trackers.
+In this work, we train $\text{Chrome}_{22}$-based classifiers on a subset of binarized response headers to evaluate their effectiveness and usefulness in web tracker detection while retaining the natural imbalance of trackers.
+We observe that response headers represent characteristics of web trackers which were previously under-explored by past research.
+They can be used to differentiate well between trackers and non-trackers and provide additional information on the nature of tracking-related entities.
+Our data-driven approach combined with binarization reveals new and confirms previously identified features.
+Applying the same methodology to request headers results in a weaker performance, thus strengthening the feasibility of using response headers for web tracker detection.
 Our two best classifiers RF and ET achieve scores ranging from 0.9 to 0.98.
@@
 However, we also observe large performance losses for $\text{Brave}_{22}$.
-The longitudinal performance of the selected ML models deteriorates only by small margins, which highlights their potential due to the changing capabilities of trackers (Fouad et al., 2018; Bujlow et al., 2017; Englehardt and Narayanan, 2016).
+The longitudinal performance of the selected classifiers deteriorates only by small margins, which highlights their potential due to the changing capabilities of trackers (Fouad et al., 2018; Bujlow et al., 2017; Englehardt and Narayanan, 2016).
 Still, periodic re-training might be necessary if the performance decreases significantly.
-We observe that response headers represent characteristics of web trackers which were previously under-explored by past research.
-They can be used to differentiate well between trackers and non-trackers and provide additional information on the nature of tracking-related entities.
-Our data-driven approach combined with binarization reveales new and confirms previously identified features.
-Applying the same methodology to request headers results in a weaker performance, thus strengthening the feasibility of using response headers for web tracker detection.
+In future research, we plan to expand on our approach of leveraging response headers for web tracker detection by considering header values in feature engineering and exploring the viability of deep-learning approaches based on response header data.
+Additionally, we intend to integrate our approach with existing solutions to develop a holistic web tracking detection system.
 Moreover, headers represent only a part of an HTTP message, and research has shown that the message bodies are also relevant when it comes to understanding and identifying trackers (Li et al., 2015; Guarino.
@@
 Emerging technologies like Large Language Models could be used to analyze bodies in an efficient and meaningful way.
-In future research, we plan to expand on our approach of leveraging response headers for web tracker detection by considering header values in feature engineering and exploring the viability of deep-learning approaches based on response header data.
-Additionally, we intend to integrate our approach with existing solutions to develop a holistic web tracking detection system.
 Acknowledgements.
-We thank our anonymous reviewers for their feedback.
+We thank our anonymous reviewers for their valuable feedback.
+Funded by the European Union – NextGenerationEU and the German Federal Ministry of Economics and Climate Protection (BMWK) in the research project ”Werk 4.0” and is supervised by the project sponsor VDI Technologiezentrum GmbH [grant number 13IK022K].
+The authors used ChatGPT4 to revise the title of this paper by asking it to make the title more interesting.
+Appendix A Availability of Code
+To strengthen the contribution and proper scientific conduct, we released all of our code and data online for everyone at:
+The datasets for the year 2022 (Raschke, 2023a) and 2023 (Raschke, 2023b) are available under the respective references.
+Appendix B Evaluation Metrics
+Here, we describe our additional metrics in more detail and their relevance for our classification problem.
+Balanced Accuracy:
+Broderson et al.
+(Brodersen et al., 2010) introduced the Balanced Accuracy (BACC) metric for imbalanced datasets in 2010 as an alternative to average accuracy scores, which are, according to the authors, problematic for CIs and imbalanced datasets.
+The metric comprises recall and true negative rate (TNR) and is defined as:
+(2) $BalancedAccuracy=\frac{1}{2}\left(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\right)$
+Log-loss Score:
+Also referred to as cross-entropy-loss, the log-loss score focuses on the predictive performance of a model.
+It describes how large the error between a prediction probability and the true label is, where a value closer to zero represents a better score.
+However, this metric can be influenced by an imbalance, i.e., the baseline log loss gets smaller with an increase in imbalance.
+Therefore, it is essential to compare the metric to its baseline value.
+For instance, the baseline for our $\text{Chrome}_{22}$ dataset with a 70:30 ratio equals 0.611  [^9] https://towardsdatascience.com/intuition-behind-log-loss-score-4e0c9979680a.
+The model’s log-loss should be lower than its baseline.
+The log-loss score is defined as:
+(3) $L_{log}(y,p)=-[y\cdot log(p)+(1-y)\cdot log(1-p)]$
+where $y$ is the label class and $p$ is the probability.
+Matthews-Correlation-Coefficient:
+The Matthews-Correla- tion-Coefficient (MCC) (Matthews, 1975), a metric that has seen increasing popularity in recent years, is defined as:
+(4) $MCC=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)\cdot(TP+FN)\cdot(TN+FP)\cdot(TN+FN)}}$
+MCC ranges from -1 to +1 and can be understood as a correlation coefficient between the true and predicted labels.
+The MCC will only achieve a high value if both class labels are predicted with high accuracy, positioning the MCC as a meaningful metric for imbalanced datasets (Boughorbel et al., 2017).
+This was further substantiated in 2020 by Chicco et al.
+(Chicco and Jurman, 2020), who outlined how accuracy and F1-Score are not reliable measures for imbalanced datasets.
+AUPRC:
+The last metric in our study is the Area Under the Precision-Recall Curve (AUPRC).
+The Precision-Recall Curve is a plot that illustrates the trade-off between precision and recall for a classifier at various thresholds.
+The AUPRC is an aggregated measure of this curve, representing the classifier’s ability to correctly identify positive instances across different thresholds.
+In imbalanced settings, where one class is significantly underrepresented, AUPRC is a more informative and reliable metric than ROC-AUC, as it focuses on the performance of the classifier on the minority class, which is of interest for us (Saito and Rehmsmeier, 2015).
+Appendix C Classifier Performance without Selected Alphabet-Services
+We performed an additional experiment to further test the robustness of our classifiers and to provide preliminary results for future research.
+The $\text{Chrome}_{22}$ classifiers were trained without HTTP responses containing Google or DoubleClick in their hostname and tested on the in-distribution dataset.
+Comparing the ET and RF with our original classifiers showed a slight decrease for AUPRC (ET: 0.959, RF: 0.962), MCC (ET: 0.876, RF: 0.872), and F1-Score (ET: 0.902, RF: 0.899).
+Important features were also similar, but the X-Frame-Options header was no longer part of the top ten features and was replaced by the Priority header.
+Testing the new classifiers on a dataset containing all of the previously excluded HTTP responses showed a larger decrease in performance.
+However, the LR model now outperformed all other classifiers – AUPRC (ET: 0.954, RF: 0.942, LR: 0.975), MCC (ET: 0.297, RF: 0.201, LR: 0.547), and F1-Score (ET: 0.652, RF: 0.559, LR: 0.866).
+Appendix D Cross-Browser and HTTP Request Classifier Performance
+This section presents our classifier performance metric results for $\text{Chrome}_{23}$, $\text{Firefox}_{22}$, $\text{Brave}_{22}$.
+In addition, it includes the results of our HTTP request header trained classifiers, including the top ten features across each classifier.
+Model Accuracy Log-Loss ROC-AUC AUPRC BACC F1-Score Precision Recall MCC FP TN FN TP
+LR 0.895 0.246 0.849 0.881 0.849 0.787 0.821 0.756 0.718 34148 561918 50425 156322
+[0.894;0.895] [0.245;0.247] [0.848;0.851] [0.88;0.882] [0.848;0.851] [0.785;0.789] [0.819;0.822] [0.754;0.759] [0.716;0.72] - - - -
+GNB 0.86 2.47 0.825 0.759 0.825 0.735 0.717 0.754 0.64 61574 534492 50925 155822
+[0.859;0.861] [2.454;2.488] [0.824;0.826] [0.757;0.76] [0.824;0.826] [0.733;0.736] [0.715;0.719] [0.752;0.756] [0.638;0.642] - - - -
+DT 0.928 1.358 0.905 0.854 0.905 0.86 0.861 0.859 0.811 28600 567466 29241 177506
+[0.927;0.929] [1.344;1.374] [0.904;0.906] [0.852;0.855] [0.904;0.906] [0.859;0.861] [0.86;0.863] [0.857;0.86] [0.81;0.813] - - - -
+RF 0.936 0.326 0.913 0.94 0.913 0.875 0.883 0.866 0.832 23712 572354 27639 179108
+[0.935;0.937] [0.32;0.333] [0.912;0.914] [0.94;0.941] [0.912;0.914] [0.874;0.876] [0.882;0.885] [0.865;0.868] [0.83;0.833] - - - -
+ET 0.944 0.381 0.92 0.936 0.92 0.888 0.905 0.873 0.851 18998 577068 26330 180417
+[0.943;0.944] [0.374;0.387] [0.92;0.921] [0.935;0.937] [0.92;0.921] [0.887;0.889] [0.904;0.906] [0.871;0.874] [0.849;0.852] - - - -
+AdaBoost 0.891 0.565 0.848 0.873 0.848 0.782 0.809 0.757 0.711 36991 559075 50140 156607
+[0.891;0.892] [0.565;0.565] [0.847;0.849] [0.872;0.875] [0.847;0.849] [0.781;0.784] [0.807;0.811] [0.756;0.76] [0.709;0.713] - - - -
+GBM 0.899 0.235 0.864 0.899 0.864 0.802 0.812 0.792 0.735 37790 558276 43056 163691
+[0.899;0.9] [0.234;0.236] [0.863;0.865] [0.898;0.9] [0.863;0.865] [0.801;0.803] [0.811;0.814] [0.789;0.794] [0.733;0.736] - - - -
+LGBM 0.927 0.177 0.899 0.939 0.899 0.856 0.871 0.841 0.807 25678 570388 32912 173835
+[0.926;0.928] [0.177;0.178] [0.898;0.9] [0.938;0.94] [0.898;0.9] [0.855;0.857] [0.87;0.873] [0.839;0.843] [0.806;0.809] - - - -
+HistGB 0.927 0.177 0.898 0.939 0.898 0.855 0.872 0.839 0.807 25504 570562 33235 173512
+[0.926;0.928] [0.176;0.178] [0.897;0.899] [0.938;0.939] [0.897;0.899] [0.854;0.857] [0.87;0.873] [0.837;0.841] [0.805;0.808] - - - -
+XGBoost 0.933 0.165 0.908 0.946 0.908 0.868 0.879 0.857 0.823 24323 571743 29621 177126
+[0.932;0.933] [0.164;0.166] [0.907;0.909] [0.946;0.947] [0.907;0.909] [0.866;0.869] [0.878;0.881] [0.854;0.858] [0.821;0.824] - - - -
+Table 6. Performance comparison of the $\text{Chrome}_{\textbf{22}}$-trained classifiers that were tested on $\text{Chrome}_{\textbf{23}}$ test data.
+Model Accuracy Log-Loss ROC-AUC AUPRC BACC F1-Score Precision Recall MCC FP TN FN TP
+LR 0.798 0.479 0.751 0.818 0.751 0.676 0.831 0.57 0.557 35881 490984 132978 176413
+[0.797;0.799] [0.477;0.48] [0.75;0.752] [0.816;0.819] [0.75;0.752] [0.675;0.678] [0.829;0.832] [0.568;0.572] [0.555;0.559] - - - -
+GNB 0.789 4.382 0.754 0.703 0.754 0.685 0.765 0.62 0.535 59045 467820 117571 191820
+[0.788;0.79] [4.359;4.401] [0.753;0.755] [0.701;0.705] [0.753;0.755] [0.683;0.686] [0.763;0.767] [0.618;0.622] [0.533;0.537] - - - -
+DT 0.861 3.225 0.826 0.814 0.826 0.786 0.907 0.694 0.699 21890 504975 94749 214642
+[0.86;0.861] [3.207;3.248] [0.825;0.827] [0.813;0.816] [0.825;0.827] [0.785;0.788] [0.906;0.908] [0.692;0.695] [0.698;0.701] - - - -
+RF 0.854 2.835 0.822 0.892 0.822 0.78 0.884 0.698 0.684 28256 498609 93560 215831
+[0.854;0.855] [2.818;2.857] [0.821;0.823] [0.828;0.83] [0.821;0.823] [0.779;0.781] [0.883;0.886] [0.696;0.699] [0.682;0.686] - - - -
+ET 0.855 2.954 0.822 0.825 0.822 0.78 0.887 0.696 0.685 27377 499488 94056 21335
+[0.854;0.856] [2.936;2.975] [0.821;0.823] [0.823;0.826] [0.821;0.823] [0.779;0.781] [0.886;0.888] [0.694;0.698] [0.683;0.687] - - - -
+AdaBoost 0.78 0.676 0.732 0.814 0.732 0.648 0.795 0.546 0.514 43497 483368 140402 168989
+[0.779;0.781] [0.676;0.676] [0.731;0.733] [0.813;0.815] [0.731;0.7333] [0.646;0.649] [0.793;0.797] [0.544;0.548] [0.512;0.516] - - - -
+GBM 0.815 0.447 0.77 0.833 0.77 0.704 0.859 0.597 0.596 30267 496598 124761 184630
+[0.814;0.815] [0.445;0.448] [0.769;0.771] [0.832;0.834] [0.769;0.771] [0.703;0.706] [0.858;0.861] [0.595;0.599] [0.594;0.598] - - - -
+LGBM 0.83 0.483 0.788 0.841 0.788 0.731 0.882 0.625 0.632 25973 500892 116153 193238
+[0.829;0.831] [0.481;0.485] [0.787;0.789] [0.84;0.842] [0.787;0.789] [0.73;0.733] [0.88;0.883] [0.623;0.627] [0.63;0.633] - - - -
+HistGB 0.83 0.478 0.788 0.843 0.788 0.731 0.881 0.625 0.631 26198 500667 115984 193407
+[0.829;0.831] [0.476;0.48] [0.787;0.789] [0.842;0.844] [0.787;0.789] [0.73;0.733] [0.879;0.882] [0.624;0.627] [0.629;0.633] - - - -
+XGBoost 0.837 0.549 0.799 0.841 0.799 0.747 0.875 0.652 0.645 28926 497939 107618 193407
+[0.836;0.838] [0.547;0.552] [0.798;0.8] [0.839;0.841] [0.798;0.8] [0.746;0.749] [0.873;0.876] [0.651;0.654] [0.643;0.647] - - - -
+Table 7. Performance comparison of the $\text{Chrome}_{\textbf{22}}$-trained classifiers that were tested on $\text{Firefox}_{\textbf{22}}$ test data.
+Model Accuracy Log-Loss ROC-AUC AUPRC BACC F1-Score Precision Recall MCC FP TN FN TP
+LR 0.947 0.174 0.728 0.06 0.728 0.091 0.05 0.506 0.147 28500 533194 1467 1502
+[0.946;0.948] [0.173;0.175] [0.719;0.738] [0.056;0.065] [0.719;0.738] [0.087;0.095] [0.048;0.052] [0.489;0.525] [0.141;0.153] - - - -
+GNB 0.896 1.847 0.825 0.064 0.825 0.071 0.037 0.754 0.152 58048 503646 731 2238
+[0.895;0.897] [1.832;1.865] [0.816;0.832] [0.06;0.067] [0.816;0.832] [0.068;0.073] [0.035;0.038] [0.736;0.767] [0.147;0.156] - - - -
+DT 0.984 0.159 0.922 0.443 0.992 0.364 0.231 0.859 0.441 8474 553220 419 2550
+[0.984;0.985] [0.154;0.164] [0.916;0.927] [0.429;0.457] [0.916;0.927] [0.354;0.372] [0.223;0.237] [0.846;0.869] [0.432;0.447] - - - -
+RF 0.985 0.063 0.923 0.728 0.923 0.383 0.247 0.86 0.456 7801 552893 415 2554
+[0.985;0.986] [0.061;0.064] [0.917;0.928] [0.713;0.745] [0.917;0.928] [0.373;0.391] [0.238;0.253] [0.848;0.87] [0.447;0.463] - - - -
+ET 0.986 0.075 0.922 0.672 0.922 0.399 0.26 0.858 0.468 7264 554430 422 2547
+[0.986;0.087] [0.073;0.078] [0.916;0.927] [0.654;0.687] [0.916;0.927] [0.387;0.407] [0.251;0.266] [0.846;0.866] [0.458;0.474] - - - -
+AdaBoost 0.94 0.669 0.734 0.042 0.734 0.085 0.046 0.525 0.142 32299 529395 1411 1558
+[0.94;0.941] [0.669;0.669] [0.725;0.742] [0.039;0.045] [0.725;0.742] [0.081;0.088] [0.044;0.048] [0.508;0.542] [0.137;0.148] - - - -
+GBM 0.963 0.158 0.835 0.439 0.835 0.167 0.095 0.706 0.25 19964 541730 873 2096
+[0.963;0.964] [0.157;0.158] [0.829;0.843] [0.418;0.46] [0.829;0.843] [0.162;0.172] [0.092;0.098] [0.693;0.72] [0.244;0.256] - - - -
+LGBM 0.979 0.109 0.86 0.537 0.86 0.267 0.163 0.74 0.341 11312 550382 773 2196
+[0.978;0.979] [0.109;0.11] [0.851;0.867] [0.518;0.559] [0.851;0.867] [0.257;0.274] [0.155;0.168] [0.724;0.755] [0.33;0.349] - - - -
+HistGB 0.979 0.109 0.861 0.549 0.861 0.267 0.163 0.742 0.341 11340 550354 766 2203
+[0.978;0.979] [0.108;0.109] [0.854;0.869] [0.531;0.57] [0.854;0.869] [0.256;0.275] [0.156;0.168] [0.727;0.758] [0.331;0.349] - - - -
+XGBoost 0.979 0.091 0.871 0.643 0.871 0.276 0.169 0.762 0.352 11148 550546 707 2262
+[0.979;0.979] [0.091;0.092] [0.864;0.878] [0.627;0.661] [0.864;0.878] [0.267;0.284] [0.162;0.174] [0.747;0.775] [0.343;0.361] - - - -
+Table 8. Performance comparison of the $\text{Chrome}_{\textbf{22}}$-trained classifiers that were tested on $\text{Brave}_{\textbf{22}}$ test data.
+Figure 9. The AdaBoost and GNB classifier improved significantly through calibration using isotonic regression.
+[…]
+[Calibrating the classifiers for potential perfomance improvements.]The AdaBoost and GNB classifier improved significantly through calibration using isotonic regression.
+Figure 10.
+The top ten features are presented in order of importance across different classifiers based on the in-distribution test set.
+The heatmap illustrates the normalized importance values for each feature and classifier.
+We calculated Permutation feature importance for LR and GNB but not for the others as we do not have any high cardinality features.
+[…]
+[Feature importance for the request-based classifiers.]The top ten features are presented in order of importance across different classifiers based on the in-distribution test set.
+The heatmap illustrates the normalized importance values for each feature and classifier.
+We calculated Permutation feature importance for LR and GNB but not for the others as we do not have any high cardinality features.
+Model Accuracy Log-Loss ROC-AUC AUPRC BACC F1-Score Precision Recall MCC FP TN FN TP
+LR 0.767 0.537 0.663 0.515 0.663 0.509 0.687 0.404 0.392 9420.0 111084.0 30559.0 20682.0
+[0.765;0.769] [0.534;0.539] [0.66;0.665] [0.511;0.52] [0.66;0.665] [0.504;0.513] [0.683;0.693] [0.4;0.408] [0.387;0.397] - - - -
+GNB 0.304 3.056 0.502 0.488 0.502 0.46 0.299 0.995 0.025 119366.0 1138.0 236.0 51005.0
+[0.302;0.305] [3.037;3.074] [0.502;0.503] [0.484;0.493] [0.502;0.503] [0.458;0.463] [0.298;0.301] [0.995;0.996] [0.021;0.029] - - - -
+DT 0.778 0.521 0.667 0.548 0.667 0.514 0.743 0.393 0.42 6965.0 113539.0 31123.0 20118.0
+[0.776;0.78] [0.518;0.523] [0.665;0.67] [0.543;0.552] [0.665;0.67] [0.509;0.518] [0.738;0.749] [0.388;0.398] [0.416;0.425] - - - -
+RF 0.778 0.52 0.667 0.548 0.667 0.514 0.743 0.393 0.42 6969.0 113535.0 31113.0 20128.0
+[0.776;0.78] [0.518;0.523] [0.665;0.67] [0.543;0.552] [0.665;0.67] [0.509;0.519] [0.738;0.749] [0.388;0.398] [0.416;0.425] - - - -
+ET 0.778 0.52 0.667 0.548 0.667 0.514 0.743 0.393 0.42 6966.0 113538.0 31123.0 20118.0
+[0.776;0.78] [0.518;0.523] [0.665;0.67] [0.543;0.552] [0.665;0.67] [0.509;0.518] [0.738;0.749] [0.388;0.398] [0.416;0.425] - - - -
+AdaBoost 0.767 0.686 0.663 0.516 0.663 0.508 0.687 0.404 0.392 9423.0 111081.0 30562.0 20679.0
+[0.765;0.769] [0.686;0.686] [0.66;0.665] [0.511;0.52] [0.66;0.665] [0.504;0.513] [0.682;0.692] [0.4;0.408] [0.387;0.397] - - - -
+GBM 0.777 0.525 0.666 0.539 0.666 0.512 0.737 0.393 0.417 7183.0 113321.0 31124.0 20117.0
+[0.775;0.779] [0.523;0.528] [0.664;0.669] [0.535;0.543] [0.664;0.669] [0.507;0.517] [0.732;0.742] [0.388;0.397] [0.412;0.422] - - - -
+LGBM 0.778 0.52 0.667 0.548 0.667 0.514 0.743 0.393 0.42 6968.0 113536.0 31123.0 20118.0
+[0.776;0.78] [0.518;0.523] [0.665;0.67] [0.543;0.552] [0.665;0.67] [0.509;0.518] [0.738;0.749] [0.388;0.398] [0.416;0.425] - - - -
+HistGB 0.778 0.52 0.667 0.548 0.667 0.514 0.743 0.393 0.42 6971.0 113533.0 31115.0 20126.0
+[0.776;0.78] [0.518;0.523] [0.665;0.67] [0.543;0.552] [0.665;0.67] [0.509;0.519] [0.738;0.749] [0.388;0.398] [0.416;0.425] - - - -
+XGBoost 0.778 0.52 0.667 0.548 0.667 0.514 0.743 0.393 0.42 6966.0 113538.0 31123.0 20118.0
+[0.776;0.78] [0.518;0.523] [0.665;0.67] [0.543;0.552] [0.665;0.67] [0.509;0.518] [0.738;0.749] [0.388;0.398] [0.416;0.425] - - - -
+Table 9. Performance comparison of the $\text{Chrome}_{\textbf{22}}$-trained classifiers with HTTP request headers that were tested on in-distribution test data.